### v1

# Primary gravitational waves at high frequencies I: Origin of suppression in the power spectrum

Abstract
The primary gravitational waves (PGWs) are generated in the early universe from the quantum vacuum during inflation. In the standard scenario of slow roll inflation, the power spectrum (PS) of PGWs over large scales, which leave the Hubble radius during inflation, is nearly scale-invariant. However, over very small scales, which never leave the Hubble radius, the PS of PGWs behaves as $k^{2}$, where $k$ denotes the wave number. In this work, we examine the PS of PGWs at such high wave numbers or frequencies when the PGWs are evolved post-inflation, through the epochs of radiation and matter domination. Firstly, we argue that the PS has to be regularized in order to truncate the unphysical $k^{2}$ rise at high frequencies. Assuming instantaneous transitions from inflation to the epochs of radiation and matter domination, we carry out the method of adiabatic regularization to arrive at the PS of PGWs over a wide range of frequencies. We show that the process of regularization truncates the $k^{2}$ rise and the PS of PGWs oscillates with a fixed amplitude about a vanishing mean value over small scales or, equivalently, at high frequencies. Secondly, we smooth the transition from inflation to radiation domination (to be precise, we smooth the ‘effective potential’ governing the equation of motion of PGWs) and examine the impact of the smoothing on the regularized PS of PGWs. With the help of a linear smoothing function, we explicitly show that the smoother transition leads to a power-law suppression in the amplitude of the oscillations (about the zero mean value) of the regularized PS of PGWs over small scales that never leave the Hubble radius during inflation. Our analysis indicates that, when transitions are involved, regularization as well as smooth transitions seem essential to ensure that the correlation functions of the PGWs in real space are well behaved. We conclude with a discussion on the different directions in which the results we have obtained need to be extended.
1 Introduction
Inflation provides an attractive mechanism to explain the origin of the primordial perturbations (see, for example, the reviews [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]). The scalar perturbations generated during inflation are primarily responsible for the anisotropies in the cosmic microwave background (CMB) [11, 12] and the large-scale structure [13, 14, 15, 16]. During inflation, in addition to the scalar perturbations, the tensor perturbations or gravitational waves (GWs) are also generated from the quantum vacuum (for the initial discussions, see, for example, Refs. [17, 18]; for recent reviews on the generation of primary and secondary GWs, see, for instance, Refs. [19, 20, 21, 22, 23]). Since the GWs interact weakly, the primary GWs (PGWs), after being generated from the quantum vacuum during inflation, propagate freely thereafter. Hence, they carry imprints of the different phases of the universe, such as the epochs of reheating, radiation and matter domination as well as any additional epochs that may have arisen in between (in this context, see, for instance, Refs. [24, 25, 26]). Therefore, observations of the power spectrum (PS) of the PGWs can, in principle, help us reconstruct the history of the early universe.
There are two important pieces of information related to inflation that we hope to glean from the direct and indirect observations of PGWs. They are the energy scale at which inflation occurred and the duration of inflation. We mentioned above that the scalar perturbations are primarily responsible for the anisotropies in the CMB. The tensor perturbations too leave specific signatures on the CMB (for a detailed discussion on the topic, see the review [27]). In particular, they are solely responsible (at least on large scales or, equivalently, at the lower multipoles) for the so-called B-mode polarization of the CMB [^1] On smaller scales or higher multipoles, gravitational lensing of the scalar perturbations can also induce the B-mode polarization in the CMB [28, 29].. The observation of the B-mode polarization in the CMB on large scales can immediately point to the energy scale of inflation. However, no observations of the B-mode polarization have yet been made over the lower multipoles of the CMB. As a result, there exists only an upper bound on the contributions of the tensor perturbations or primordial GWs to the anisotropies in the CMB (see Refs. [11, 12]; for the prospects of detection of the B-mode in the near future, see Ref. [30]). The latest CMB observations constrain the dimensionless tensor-to-scalar ratio $r$ to be $r<0.036$ [31]. This roughly translates to the upper bound on the Hubble scale during inflation, say, $H_{{}_{\mathrm{I}}}$, to be $H_{{}_{\mathrm{I}}}/M_{{}_{\mathrm{Pl}}}\simeq 10^{-5}$, where $M_{{}_{\mathrm{Pl}}}$ denotes the Planck mass.
The constraints from the CMB on the primordial GWs are indirect. We should mention that the frequency, say, $f$, of the primordial GWs is related to the wave number $k$ through the relation
$f\equiv\frac{k/a_{0}}{2\pi}=1.55\times 10^{-15}\left(\frac{k/a_{0}}{1\,\mathrm{Mpc}^{-1}}\right)\,\mathrm{Hz},$ (1.1)
where $a_{0}$ is the scale factor of the Friedmann-Lemaître-Robertson-Walker (FLRW) universe today. As convenient, we shall describe the PS of GWs either in terms of wave numbers or frequencies. Note that the CMB constrains the primordial GWs only over large scales corresponding to wave numbers $10^{-4}\lesssim k/a_{0}\lesssim 1\,\rm{Mpc}^{-1}$ or, equivalently, over very small frequencies in the range of $10^{-19}\lesssim f\lesssim 10^{-15}\,\rm{Hz}$. During the past decade, the direct detection of GWs from astrophysical sources, such as coalescing compact binary black holes and neutron stars, by the Ligo-Virgo-Kagra (LVK) collaboration has opened a new window to the universe (in this context, see, for instance, Refs. [32, 33, 34, 35, 36, 37]). The GWs generated in the primordial universe are expected to be stochastic in nature. The detection of GWs from astrophysical sources by the LVK collaboration has led to the expectation that the direct detection of a stochastic GW background (SGWB) may be around the corner. But, despite a few rounds of observations, the LVK collaboration has not yet detected a SGWB over the range of frequencies (say, $1\lesssim f\lesssim 10^{2}\,\mathrm{Hz}$) that its detectors are sensitive to [38, 39, 40]. However, about two years ago, the Pulsar Timing Arrays (PTAs)—such as NANOGrav [41, 42], EPTA (including the data from InPTA) [43, 44], PPTA [45, 46], and CPTA [47]—reported the detection of a SGWB in the range of frequencies $10^{-9}\lesssim f\lesssim 10^{-6}\,\mathrm{Hz}$ [^2] We should mention that the detection of the SGWB by the PTAs also constitute an indirect detection (for a discussion in this context, see, for instance, Ref. [48]).. Many ongoing and forthcoming GW observatories, such as the Square Kilometer Array (SKA) ($10^{-9}\lesssim f\lesssim 10^{-6}\,\mathrm{Hz}$) [49], Laser Interferometer Space Antenna (LISA) ($10^{-4}\lesssim f\lesssim 0.1\,\rm{Hz}$) [50, 51], TianQin and Taiji ($10^{-4}\lesssim f\lesssim 0.1\,\rm{Hz}$) [52, 53, 54], Big Bang Observer (BBO) ($0.1\lesssim f\lesssim 1\,\rm{Hz}$) [55, 56, 57, 58], Decihertz Interferometer Gravitational wave Observatory (DECIGO) ($0.1\lesssim f\lesssim 10\,\rm{Hz}$) [59, 60], Matter-wave Atomic Gradiometer Interferometric Sensor (MAGIS) ($0.1\lesssim f\lesssim 10\,\rm{Hz}$) [61], Cosmic Explorer (CE) ($1\lesssim f\lesssim 10^{2}\,\rm{Hz}$) [62, 63] and Einstein Telescope (ET) ($1\lesssim f\lesssim 10^{3}\,\rm{Hz}$) [64, 65, 66], are expected to operate over the indicated range of frequencies. At high frequencies ($10^{3}\lesssim f\lesssim 10^{20}\,\rm{Hz}$), we already have some upper bounds on the characteristic strain of GWs from Bulk Acoustic Wave devices (BAW) [67, 68], Absolute Radiometer for Cosmology, Astrophysics and Diffuse Emission (ARCADE) [69], Optical Search for QED Vacuum Birefringence, Axions and Photon Regeneration (OSQAR) [70], and the CERN Axion Solar Telescope (CAST) [71]. In addition, a set of detectors have been proposed to observe GWs at these high frequencies (for broad discussions in this context, see, Refs. [72, 73, 74, 75, 76]). Specific detectors include Axion Dark Matter eXperiment (ADMX) [77], Superconducting Quantum Materials and Systems Center (SQMS) [78], International Axion Observatory-heterodyne+single photon detectors (IAXO-HET+SPD) [79], electromagnetic Gaussian beams (GB) [80], Joint Undertaking on the Research for Axion-like particles (JURA) [81], and International Axion Observatory (IAXO) [82]. During the coming decade or two, these observatories are anticipated to provide further constraints on the amplitude and shape of the PS of primordial GWs.
In this work, we shall be interested in understanding the behavior of the PS of PGWs over small scales or, equivalently, at high frequencies. Specifically, we shall be interested in wave numbers that never leave the Hubble radius during the course of cosmological evolution. Let $k_{\mathrm{e}}$ denote the wave number that leaves the Hubble radius at the end of inflation. We shall now briefly discuss the expected behavior of the PS of PGWs over small scales such that $k\gtrsim k_{\mathrm{e}}$. Since these wave numbers remain inside the Hubble radius from the early stages of inflation until today, the mode functions describing the PGWs associated with these wave numbers will always behave in the same manner (barring an overall time-dependent function that depends inversely on the scale factor, say, $a$) as the mode functions in Minkowski spacetime. Then, it can be immediately shown that, over such wave numbers, at any time, the PS of PGWs will behave as $k^{2}$. Moreover, as we shall illustrate in due course, when the PS of PGWs is evaluated today, the $k^{2}$ rise will suggest that their strengths at high frequencies can become comparable to the sensitivities of the ongoing and forthcoming GW observatories (in this context, see Fig. 3). It has been argued that the PS of PGWs have to be regularized, which can truncate the $k^{2}$ rise in these quantities at large wave numbers (see Ref. [83]; for related discussions on the scalar power spectrum, see Refs. [84, 85, 86, 87, 88, 89, 90]; in this context, also see Ref. [91]). As we shall discuss, in the context of PGWs, the process of regularization is essential to avoid the $k^{2}$ rise in the PS on small scales.
There is another aspect of the background cosmological dynamics that is expected to modify the PS of PGWs around the wave number $k_{\mathrm{e}}$. It has been suggested in the literature (we should say with inadequate proof) that the PS of GWs will exhibit an exponential suppression over wave numbers $k\gtrsim k_{\mathrm{e}}$ (see, for instance, Refs. [92, 93, 94, 23]). Barring an occasional exception, the PGWs are typically evolved from their origin during inflation to the present epoch assuming that the transitions from one epoch to another (say, from inflation to reheating or the epoch of radiation domination) are instantaneous or abrupt. Realistically, these transitions should be infinitely smooth and recent arguments in the literature suggest that smoothing the transitions can lead to a sharp cut-off in the PS of PGWs for $k\gtrsim k_{\mathrm{e}}$ (in this context, see Refs. [90, 95]).
Our aim in this work is to carefully examine the effects due to regularization as well as smoothing of the transitions on the PS of PGWs. To understand the behavior around $k_{\mathrm{e}}$, we use the method of adiabatic regularization to calculate the observed PS of PGWs today. We also smooth the transition from inflation to the epoch of radiation domination in a specific manner and explicitly evaluate the corresponding modifications to the PS of PGWs. We show that the regularization leads to a PS which oscillates with a fixed amplitude about a vanishing mean value over $k\gtrsim k_{\mathrm{e}}$. We further illustrate that, at such high wave numbers, the smoother transition leads to a power-law suppression of the amplitude of the oscillations in the regularized PS of PGWs.
This paper is organized as follows. In the following section, we shall describe some of the essential properties of PGWs. We shall describe the equation of motion satisfied by the PGWs, their quantization, and introduce the PS associated with them. We shall also discuss the generic form of the PS expected in the post-inflationary universe. In Sec. 3, we shall discuss the popular case of instantaneous transitions from de Sitter inflation to the epochs of radiation and matter domination, and evaluate the resulting PS of PGWs. In Sec. 4, we shall initially discuss the need for the regularization of the PS of PGWs. We shall also describe the manner in which the PS of PGWs can be regularized using the method of adiabatic subtraction. Moreover, in these sections, we shall clarify a few points related to the amplitude and shape of the actual as well as the regularized PS of PGWs. In particular, we shall show that the regularized PS of PGWs, evaluated during the epochs of radiation and matter domination, oscillates with a constant amplitude about zero over $k\gtrsim k_{\mathrm{e}}$. In Sec. 5, we shall smooth the transition from power-law inflation (in fact, we shall smooth the ‘effective potential’ that governs the dynamics of PGWs) to the epoch of radiation domination with the aid of a linear function and determine the scale factor during the transition. In Sec. 6, we shall evolve the mode functions of the PGWs across the smooth linear transition and illustrate that, over $k\gtrsim k_{\mathrm{e}}$, the regularized PS of PGWs contains oscillations about a vanishing mean and that the amplitude of the oscillations exhibits a power-law suppression. Finally, in Sec. 7, after a quick summary of the results we have obtained, we shall outline a few related issues that need to be examined further.
At this point in our discussion, we should make a few clarifying remarks concerning the conventions and notations that we shall work with. We shall work with natural units such that $\hbar=c=1$ and set the reduced Planck mass to be $M_{{}_{\mathrm{Pl}}}=\left(8\pi G\right)^{-1/2}$. We shall adopt the signature of the metric to be (,+,+,+)(-,+,+,+). Note that Latin indices shall represent the spatial coordinates, except for $k$ which shall be reserved for denoting the wave number. We shall assume the background to be the spatially flat, FLRW universe described by the following line-element:
${\mathrm{d}}s^{2}=-{\mathrm{d}}t^{2}+a^{2}(t){\mathrm{d}}{\bm{x}}^{2}=a^{2}(\eta)\left(-{\mathrm{d}}\eta^{2}+{\mathrm{d}}{\bm{x}}^{2}\right),$ (1.2)
where $t$ and $\eta$ denote the cosmic and conformal time coordinates, and $a$ denotes the scale factor. The overdots and overprimes shall denote differentiation with respect to the cosmic and conformal times, respectively.
2 Essentials of PGWs
In this section, we shall describe aspects of PGWs that are essential for our discussion. We shall first introduce the equation of motion governing PGWs. We shall also describe the quantization of PGWs and discuss the associated PS.
2.1 Equation of motion of PGWs and quantization
Let $h_{ij}$ denote the tensor perturbations characterizing the GWs in a FLRW universe. When these tensor perturbations are taken into account, the line-element (1.2) describing the spatially flat, FLRW universe is modified to be (see, for instance, Refs. [1, 96])
${\mathrm{d}}s^{2}$ $=-{\mathrm{d}}t^{2}+a^{2}(t)\left[\delta_{ij}+h_{ij}(t,{\bm{x}})\right]{\mathrm{d}}x^{i}{\mathrm{d}}x^{j}$
$=a^{2}(\eta)\left\{-{\mathrm{d}}\eta^{2}+\left[\delta_{ij}+h_{ij}(\eta,{\bm{x}})\right]{\mathrm{d}}x^{i}{\mathrm{d}}x^{j}\right\}.$ (2.1)
Since the tensor perturbations $h_{ij}$ are transverse and traceless, they satisfy the conditions
$\delta^{ij}\partial_{i}h_{jl}=0,\;\delta^{ij}h_{ij}=0.$ (2.2)
Also, in our discussion, we shall assume that no sources with an anisotropic stress-energy tensor are present. Under these conditions, the Einstein’s equations at the first order in the perturbations lead to the following equation of motion for the PGWs (for a derivation, see App. A):
$h_{ij}^{\prime\prime}+2{\cal H}h_{ij}^{\prime}-\partial^{2}h_{ij}=0,$ (2.3)
where ${\cal H}=a^{\prime}/a=aH$ is the conformal Hubble parameter, with $H=\dot{a}/a$ being the Hubble parameter, and $\partial^{2}=\delta^{ij}\partial_{i}\partial_{j}$.
As we mentioned earlier, the PGWs originate from the quantum vacuum. On quantization, the tensor perturbations $h_{ij}$ can be decomposed in terms of the rescaled mode functions, say, $\mu_{k}(\eta)$, as follows [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]:
$\hat{h}_{ij}(\eta,{\bm{x}})=\frac{\sqrt{2}}{M_{{}_{\mathrm{Pl}}}a(\eta)}\sum_{\lambda={(+,\times)}}\int\frac{{\mathrm{d}}^{3}{\bm{k}}}{(2\pi)^{3/2}}\varepsilon^{\lambda}_{ij}(\hat{\bm{n}})\left[\hat{a}^{\lambda}_{\bm{k}}\mu_{k}(\eta)\mathrm{e}^{i{\bm{k}}\cdot{\bm{x}}}+\hat{a}^{\lambda{\dagger}}_{\bm{k}}\mu^{\ast}_{k}(\eta)\mathrm{e}^{-i{\bm{k}}\cdot{\bm{x}}}\right],$ (2.4)
[…]
$\delta^{ij}k_{i}\varepsilon_{jl}^{\lambda}(\hat{\bm{n}})=0,\;\delta^{ij}\varepsilon_{ij}^{\lambda}(\hat{\bm{n}})=0.$ (2.6)
[…]
$\delta^{im}\delta^{jn}\varepsilon_{mn}^{\lambda}(\hat{\bm{n}})\varepsilon_{ij}^{\lambda^{\prime}}(\hat{\bm{n}})=2\,\delta^{\lambda\lambda^{\prime}}.$ (2.7)
In the Fourier decomposition (2.4) of the tensor perturbations, the operators $\hat{a}_{\bm{k}}^{\lambda}$ and $\hat{a}^{\lambda\dagger}_{\bm{k}}$ denote the annihilation and creation operators corresponding to the tensor modes associated with the wave vector $\bm{k}$} and polarization $\lambda$. The annihilation and creation operators obey the standard commutation relations, viz.
$\left[\hat{a}^{\lambda}_{\bm{k}},\hat{a}_{\bm{k}^{\prime}}^{\lambda^{\prime}}\right]=\left[\hat{a}^{\lambda{\dagger}}_{\bm{k}},\hat{a}_{\bm{k}^{\prime}}^{\lambda^{\prime}\dagger}\right]=0,\quad\left[\hat{a}^{\lambda}_{\bm{k}},\hat{a}_{\bm{k}^{\prime}}^{\lambda^{\prime}\dagger}\right]=\delta^{(3)}\left({\bm{k}}-{\bm{k}}^{\prime}\right)\delta^{\lambda\lambda^{\prime}}.$ (2.8)
In the absence of sources with anisotropic stresses, the equation of motion (2.3) for $h_{ij}$ leads to the following differential equation for the rescaled mode function $\mu_{k}(\eta)$ [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]:
$\mu_{k}^{\prime\prime}+\left(k^{2}-\frac{a^{\prime\prime}}{a}\right)\mu_{k}=0.$ (2.9)
Clearly, in terms of the variable $\mu_{k}(\eta)$, the tensor mode functions correspond to solutions of a time-dependent oscillator whose frequency is determined by the behavior of the scale factor.
2.2 PS of PGWs
In this work, we shall focus on the PS of the primordial tensor perturbations, which is one of the quantities of observational interest. The PS is a measure of the two-point correlation of the tensor perturbations in Fourier space. The PS of PGWs at the conformal time $\eta$, say, $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$, is defined through the relation [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]
$\delta^{im}\delta^{jn}\left\langle 0\left|\hat{h}_{mn}(\eta,{\bm{x}})\hat{h}_{ij}(\eta,{\bm{x}})\right|0\right\rangle=\int_{0}^{\infty}\frac{\mathrm{d}k}{k}\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta),$ (2.10)
where, as is usually done, we have assumed that the expectation value on the left hand side is evaluated in the quantum vacuum. The vacuum state $|0\rangle$ is defined as ${\hat{a}}_{\bm{k}}^{\lambda}|0\rangle=0$ for all $\bm{k}$} and $\lambda$. On utilizing the decomposition (2.4) of the quantized tensor perturbations $\hat{h}_{ij}$ in terms of the rescaled mode functions $\mu_{k}(\eta)$, the PS of PGWs at any time can be expressed as
$\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)=\frac{8}{M_{{}_{\mathrm{Pl}}}^{2}}\frac{k^{3}}{2\pi^{2}}\frac{|\mu_{k}(\eta)|^{2}}{a^{2}(\eta)}.$ (2.11)
2.3 PGWs in the post-inflationary universe
Evidently, to arrive at the PS of PGWs today, we need to evolve the tensor perturbations in the post-inflationary universe. After inflation, the different epochs of the universe can be described by an equation-of-state parameter, say, $w$, that is a constant and is greater than 1/3-1/3. In such situations, the scale factor describing the spatially flat, FLRW universe can be written as
$a(\eta)=a_{w}\left(\eta-\eta_{w}\right)^{2/(1+3w)}.$ (2.12)
The two constants $a_{w}$ and $\eta_{w}$ are to be determined by matching the scale factor and its time derivative at the onset of the epoch, with those of the previous epoch. For the above $a(\eta)$, we have
$\frac{a^{\prime\prime}}{a}=\frac{2(1-3w)}{(1+3w)^{2}(\eta-\eta_{w})^{2}}.$ (2.13)
In such a case, the general solution to the rescaled mode function $\mu_{k}(\eta)$ that describes the PGWs [cf. Eq. (2.9)] can be written as
$\mu_{k}(\eta)=\alpha_{k}^{w}m_{k}(\eta)+\beta_{k}^{w}n_{k}(\eta),$ (2.14)
where $(\alpha_{k}^{w},\beta_{k}^{w})$ are the so-called Bogoliubov coefficients. As in the case of the constants describing the scale factor (2.12), the coefficients $(\alpha_{k}^{w},\beta_{k}^{w})$ are to be determined by the boundary conditions on the rescaled mode function $\mu_{k}(\eta)$ imposed at the beginning of the epoch. The functions $m_{k}(\eta)$ and $n_{k}(\eta)$ are given by
$m_{k}(\eta)$ $=x^{1/2}\mathrm{H}_{\nu(w)}^{(2)}(x),$ (2.15a)
$n_{k}(\eta)$ $=x^{1/2}\mathrm{H}_{\nu(w)}^{(1)}(x)=m_{k}^{*}(\eta),$ (2.15b)
where $x=k(\eta-\eta_{w})$ and $\nu(w)=3(1-w)/[2(1+3w)]$, while $\mathrm{H}_{\nu}^{(1)}(x)$ and $\mathrm{H}_{\nu}^{(2)}(x)$ are the Hankel functions of types one and two, respectively, and of order $\nu$.
During the era where the universe is dominated by radiation, $w=1/3$ and $\nu=1/2$. In such a case, $a^{\prime\prime}/a=0$, and the function $m_{k}(\eta)$ has the following simple form [99]:
$m_{k}(\eta)=i\sqrt{\frac{2}{\pi}}\mathrm{e}^{-ix}=n_{k}^{\ast}(\eta).$ (2.16)
On substituting the corresponding $\mu_{k}(\eta)$ in the expression (2.11), we obtain the PS of PGWs during the radiation-dominated epoch to be
[…]
This formula is general and it has a distinct advantage. To predict $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$, all we need to do is determine the form of the coefficients $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ and substitute them in the above expression.
As we indicated above, the coefficients $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ are determined by the boundary conditions imposed on the rescaled mode function $\mu_{k}(\eta)$ at the beginning of the epoch of radiation domination. In other words, the coefficients depend on the history of the universe prior to the era of radiation domination. Let us now consider the special case wherein the universe, before the radiation-dominated era, underwent a phase of cosmic inflation. During inflation, the scale factor is typically of the form
$a(\eta)=\ell_{0}(-\eta)^{q},$ (2.18)
where q<1q<-1 in power-law inflation and $q=-1-\epsilon_{1}$ (with $\epsilon_{1}=-\dot{H}/H^{2}\ll 1$ being the first slow-roll parameter) in slow-roll inflation. Also, $\ell_{0}$ determines the scale of inflation. Moreover, note that exact de Sitter inflation corresponds to q=1q=-1. Further, in the inflationary scenario, it is common to impose the so-called Bunch-Davies initial conditions on the perturbations when the modes are in the sub-Hubble domain during the early stages [100]. This corresponds to assuming that the tensor perturbations are in the vacuum state. Under such assumptions, the solution for the rescaled mode function $\mu_{k}(\eta)$ during inflation can be written as
$\mu_{k}(\eta)=A_{k}f_{k}(y),$ (2.19)
where $y\equiv-k\eta>0$, and the function $f_{k}(y)$ is given by
$f_{k}(\eta)=y^{1/2}\,\mathrm{H}_{\mu}^{(1)}(y),$ (2.20)
with $\mu=1/2-q$. The constant $A_{k}$ is given by
$A_{k}=\sqrt{\frac{\pi}{2}}\frac{1}{\sqrt{2k}}\mathrm{e}^{-i\pi q/2+i\pi/2+ik\eta_{\mathrm{i}}}=\sqrt{\frac{\pi}{2}}\frac{1}{\sqrt{2k}}\mathrm{e}^{i\pi\mu/2+i\pi/4+ik\eta_{\mathrm{i}}},$ (2.21)
where $k\eta_{\mathrm{i}}$ is an arbitrary phase factor. We should clarify that inflation could have occurred either immediately before the epoch of radiation domination or there could have been other phases between inflation and the radiation-dominated era, such as the phase of reheating. Whatever the details of the scenario that one considers, the coefficients $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ will always be proportional to $A_{k}$, and they will satisfy the normalization condition
$|\alpha_{k}^{\mathrm{r}}|^{2}-|\beta_{k}^{\mathrm{r}}|^{2}=|A_{k}|^{2}.$ (2.22)
Therefore, the power spectrum (2.17) can be expressed as
[…]
and, in this expression, we have set
$\bar{y}_{\mathrm{e}}=\frac{k}{k_{\mathrm{e}}}$ (2.24)
where $k_{\mathrm{e}}=a_{\mathrm{e}}H_{\mathrm{e}}$, with $a_{\mathrm{e}}$ and $H_{\mathrm{e}}$ denoting the scale factor and the Hubble parameter at the end of inflation. We should emphasize again that the above result is very general and, to arrive at explicit expressions for the PS of PGWs, we need to evaluate the coefficients $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$. In order to calculate these coefficients, we need to specify the manner in which inflation and the epoch of radiation domination are related.
3 Instantaneous transition from de Sitter inflation to radiation and matter domination
In this section, we shall consider the simple case of instantaneous transition from de Sitter inflation to the epoch of radiation domination and calculate the coefficients $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$. Thereafter, assuming instantaneous transition from radiation to matter domination, we shall further evolve the mode functions describing the PGWs through the epoch of matter domination. We shall also take the opportunity to highlight a few points related to the resulting PS of PGWs evaluated during the epochs of radiation and matter domination.
3.1 PS during radiation domination
We shall now discuss the simplest of scenarios which allows us to calculate the coefficients $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ easily. We shall consider the popular model wherein de Sitter inflation ends, say, at the conformal time $\eta_{\mathrm{e}}$, and is immediately followed by the epoch of radiation domination. In this scenario, the transition between inflation and the radiation dominated epoch is sharp and reheating is instantaneous. Recall that, during de Sitter inflation and radiation domination, the scale factors are given by [cf. Eqs. (2.18) and (2.12)]
$a(\eta)$ $=-\frac{\ell_{0}}{\eta},$ (3.1a)
$a(\eta)$ $=a_{\mathrm{r}}(\eta-\eta_{\mathrm{r}}).$ (3.1b)
[…]
$\alpha_{k}^{\mathrm{r}}$ $=\frac{A_{k}}{W_{k}^{\mathrm{r}}}\left[f_{k}(y_{\mathrm{e}})n_{k}^{\prime}(x_{\mathrm{e}})-f_{k}^{\prime}(y_{\mathrm{e}})n_{k}(x_{\mathrm{e}})\right],$ (3.2a)
$\beta_{k}^{\mathrm{r}}$ $=-\frac{A_{k}}{W_{k}^{\mathrm{r}}}\left[f_{k}(y_{\mathrm{e}})m_{k}^{\prime}(x_{\mathrm{e}})-f_{k}^{\prime}(y_{\mathrm{e}})m_{k}(x_{\mathrm{e}})\right],$ (3.2b)
where $x_{\mathrm{e}}\equiv k(\eta_{\mathrm{e}}-\eta_{\mathrm{r}})=-k\eta_{\mathrm{e}}\equiv y_{\mathrm{e}}=k/k_{\mathrm{e}}=\bar{y}_{\mathrm{e}}$ and $k_{\mathrm{e}}=-1/\eta_{\mathrm{e}}$. The quantity $W_{k}^{\mathrm{r}}$ is the Wronskian defined as
$W_{k}^{\mathrm{r}}(\eta)\equiv m_{k}(\eta)n_{k}^{\prime}(\eta)-m_{k}^{\prime}(\eta)n_{k}(\eta).$ (3.3)
The equation of motion for $\mu_{k}(\eta)$ implies that the Wronskian is a constant and, using the forms of $m_{k}(\eta)$ and $n_{k}(\eta)$ in the radiation-dominated epoch [cf. Eq. (2.16)], it can be established that $W_{k}^{\mathrm{r}}=4ik/\pi$. Then, straightforward manipulations lead to
$\alpha_{k}^{\mathrm{r}}$ $=-\frac{iA_{k}}{2y_{\mathrm{e}}^{2}}(1-2iy_{\mathrm{e}}-2y_{\mathrm{e}}^{2})\mathrm{e}^{2iy_{\mathrm{e}}},$ (3.4a)
$\beta_{k}^{\mathrm{r}}$ $=-\frac{iA_{k}}{2y_{\mathrm{e}}^{2}},$ (3.4b)
and it can be easily verified that $|\alpha_{k}^{\mathrm{r}}|^{2}-|\beta_{k}^{\mathrm{r}}|^{2}=|A_{k}|^{2}$, as required. On substituting the coefficients $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ in the expression (2.3) for the PS of PGWs during radiation domination, we obtain that
$\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ $=\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\frac{1}{y_{\mathrm{e}}^{2}}\biggl\{y_{\mathrm{e}}^{4}+\frac{1}{2}-\frac{1}{2}\left(1-2y_{\mathrm{e}}^{2}\right)\cos\left[2(x-y_{\mathrm{e}})\right]$
$\quad+y_{\mathrm{e}}\sin\left[2(x-y_{\mathrm{e}})\right]\biggr\},$ (3.5)
where, recall that, $x=k(\eta-\eta_{\mathrm{r}})=ay_{\mathrm{e}}/a_{\mathrm{e}}$.
The above PS is valid at any time during the epoch of radiation domination. Let us now discuss in detail the shape of the PS when it is evaluated at the time of radiation-matter equality, say, $\eta_{\mathrm{eq}}$. Note that $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta_{\mathrm{eq}})$ is controlled by two quantities, $y_{\mathrm{e}}=k/k_{\mathrm{e}}$, where $k_{\mathrm{e}}$ is the wave number that leaves the Hubble radius at the end of inflation, and $x_{\mathrm{eq}}=k/k_{\mathrm{eq}}$, where $k_{\mathrm{eq}}=a_{\mathrm{eq}}H_{\mathrm{eq}}$ is the wave number that re-enters the Hubble radius at the time of radiation-matter equality. For wave numbers that are on super-Hubble scales at the end of inflation $y_{\mathrm{e}}\ll 1$, and in such a limit, we find that the PS given by Eq. (3.5) reduces to
$\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)\simeq\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\frac{1}{y_{\mathrm{e}}^{2}}\sin^{2}(x).$ (3.6)
In this domain, i.e. over $y_{\mathrm{e}}\ll 1$, let us focus on wave numbers that have remained in the super-Hubble domain in the subsequent radiation-dominated era. This corresponds to considering the limit $x\ll 1$. At the time of radiation-matter equality, such a limit corresponds to $x_{\mathrm{eq}}\ll 1$ or, equivalently, $k\ll k_{\mathrm{eq}}$. In such a case, we obtain the PS to be
$\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)=\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\frac{x^{2}}{y_{\mathrm{e}}^{2}}=\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2},$ (3.7)
since $x=ay_{\mathrm{e}}/a_{\mathrm{e}}$. In other words, the PS of PGWs is scale-invariant in this regime. For modes that were in the super-Hubble domain at the end of inflation but have then re-entered the Hubble radius during the radiation-dominated epoch, we have $x\gg 1$. At the time of radiation-matter equality, such a limit corresponds to $k_{\mathrm{eq}}\ll k<k_{\mathrm{e}}$. In such a case, the PS is still given by Eq. (3.6) above. It should be clear that, over this regime, the PS behaves as $k^{-2}$ with superimposed oscillations. Finally, over wave numbers that are in the sub-Hubble regime at the end of inflation, i.e. $y_{\mathrm{e}}\gg 1$ or $k\gg k_{\mathrm{e}}$, and, therefore, always remain in the sub-Hubble domain subsequently, we find that the PS behaves as
$\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)\simeq\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}y_{\mathrm{e}}^{2}.$ (3.8)
In other words, the PS of PGWs rises as $k^{2}$ over wave numbers such that $k\gg k_{\mathrm{e}}$.
[…]
$\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a_{\mathrm{eq}}}\right)^{2}$ $=\frac{\pi}{\sqrt{90}}\left(\frac{H_{0}/h}{M_{{}_{\mathrm{Pl}}}}\right)\left(\frac{g_{s,\mathrm{rh}}}{g_{s,\mathrm{eq}}}\right)^{-2/3}\left(\frac{g_{\ast,\mathrm{rh}}}{g_{\ast,\mathrm{eq}}}\right)^{1/2}\,\left(\frac{g_{\ast,\mathrm{rh}}^{1/4}T_{\mathrm{rh}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\frac{\left(\Omega_{\mathrm{c}}h^{2}+\Omega_{\mathrm{b}}h^{2}\right)^{2}}{\left(\Omega_{\gamma}h^{2}+\Omega_{\nu}h^{2}\right)^{3/2}}$
$\simeq 3.725\times 10^{-75}\left(\frac{g_{*,\mathrm{rh}}^{1/4}T_{\mathrm{rh}}}{10^{9}\,\mathrm{GeV}}\right)^{2},$ (3.9)
where $a_{\mathrm{eq}}$ denotes the scale factor at the time of radiation-matter equality. Also, $T_{\mathrm{rh}}$ represents the reheating temperature and $H_{0}=2.133\,h\times 10^{-42}\,\mathrm{GeV}$ is the value of the Hubble parameter today. Moreover, $(g_{\ast,\mathrm{rh}},g_{\ast,\mathrm{eq}})$ and $(g_{s,\mathrm{rh}},g_{s,\mathrm{eq}})$ denote the number of effective relativistic degrees of freedom that contribute to the energy density of radiation and entropy at the time of reheating and radiation-matter equality, respectively. The quantities $\Omega_{\mathrm{c}}\,h^{2}=0.12011$, $\Omega_{\mathrm{b}}\,h^{2}=0.0223$, $\Omega_{\gamma}\,h^{2}=2.47\times 10^{-5}$, $\Omega_{\nu}\,h^{2}=1.68\times 10^{-5}$ are the present day values for the dimensionless densities of cold dark matter, baryons, photons and neutrinos, respectively [101]. Further, to evaluate the argument of the trigonometric functions appearing in the PS, we need to evaluate the quantity $x$. Recall that
$x=k(\eta-\eta_{\mathrm{r}})=\frac{k}{aH}=\frac{k}{a_{\mathrm{e}}H_{\mathrm{e}}}\frac{a_{\mathrm{e}}H_{\mathrm{e}}}{aH}=y_{\mathrm{e}}\frac{a}{a_{\mathrm{e}}},$ (3.10)
where we have used the fact that $aH={\cal H}=(\eta-\eta_{\mathrm{r}})^{-1}$ in the radiation-dominated era. Therefore, in order to calculate $x_{\mathrm{eq}}$, we need to evaluate the ratio, $a_{\mathrm{eq}}/a_{\mathrm{e}}$ which can be expressed as
$\frac{a_{\mathrm{eq}}}{a_{\mathrm{e}}}$ $=\left(\frac{\pi^{2}}{90}\right)^{1/4}\left(\frac{H_{0}/h}{M_{{}_{\mathrm{Pl}}}}\right)^{-1/2}\left(\frac{g_{s,\mathrm{rh}}}{g_{s,\mathrm{eq}}}\right)^{1/3}\left(\frac{g_{\ast,\mathrm{eq}}}{g_{\ast,\mathrm{rh}}}\right)^{1/4}\,\left(\frac{g_{\ast,\mathrm{rh}}^{1/4}T_{\mathrm{rh}}}{M_{{}_{\mathrm{Pl}}}}\right)\frac{(\Omega_{\gamma}h^{2}+\Omega_{\nu}h^{2})^{3/4}}{(\Omega_{\mathrm{c}}h^{2}+\Omega_{\mathrm{b}}h^{2})}$
$\simeq 9.188\times 10^{17}\left(\frac{g_{\ast,\mathrm{rh}}^{1/4}T_{\mathrm{rh}}}{10^{9}\,\mathrm{GeV}}\right).$ (3.11)
Using these results, in Fig. 1, we have plotted the PS of PGWs evaluated at the time of radiation-matter equality as a function of $y_{\mathrm{e}}$. The figure clearly reflects the behavior of the PS in the three different limits we discussed above [as given by Eqs. (3.6), (3.7) and (3.8))].
[…]
3.2 PS during matter domination
Let us now arrive at the PS during the epoch of matter domination that follows the epoch of radiation domination. In order to do so, we need to evolve the rescaled mode functions $\mu_{k}(\eta)$ post the epoch of radiation-matter equality. During matter domination, $w=0$ and hence the scale factor can be expressed as [cf. Eq. (2.12)]
$a(\eta)=a_{\mathrm{m}}(\eta-\eta_{\mathrm{m}})^{2},$ (3.12)
where $a_{\mathrm{m}}$ and $\eta_{\mathrm{m}}$ are constants. We shall assume that the transition from radiation to matter domination is instantaneous. On matching the above scale factor and its time derivative at $\eta_{\mathrm{eq}}$ with the values from the earlier epoch of radiation domination [cf. Eq. (3.1b)], we obtain that
$a_{\mathrm{m}}$ $=\frac{a_{\mathrm{r}}}{2(\eta_{\mathrm{eq}}-\eta_{\mathrm{m}})},$ (3.13a)
$\eta_{\mathrm{m}}$ $=-\eta_{\mathrm{eq}}+2\eta_{\mathrm{r}}=-\eta_{\mathrm{eq}}+4\eta_{\mathrm{e}}.$ (3.13b)
During the epoch of matter domination, the general solution for the rescaled mode function $\mu_{k}(\eta)$ can be expressed as [cf. Eq. (2.15)]
$\mu_{k}(\eta)=\alpha_{k}^{\mathrm{m}}p_{k}(\eta)+\beta_{k}^{\mathrm{m}}q_{k}(\eta),$ (3.14)
where the functions $p_{k}(\eta)$ and $q_{k}(\eta)$ are given by
$p_{k}(\eta)$ $=z^{1/2}H_{3/2}^{(2)}(z)=\sqrt{\frac{2}{\pi}}\mathrm{e}^{-iz}\left(-1+\frac{i}{z}\right)=q_{k}^{\ast},$ (3.15)
with $z=k(\eta-\eta_{\mathrm{m}})$. On substituting the above mode function in the expression (2.11), we find that the PS of PGWs during the epoch of matter domination can be written as
[…]
As in the case of the Bogoliubov coefficients $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ [cf. Eq. (2.22)], it can be established that the coefficients $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$ satisfy the condition
$|\alpha_{k}^{\mathrm{m}}|^{2}-|\beta_{k}^{\mathrm{m}}|^{2}=|A_{k}|^{2}.$ (3.17)
Using this Wronskian condition, we can rewrite the PS above as
$\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ $=\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\bar{y}_{\mathrm{e}}^{2}\Biggl\{\left(1+2\,\left|\frac{\beta_{k}^{\mathrm{m}}}{A_{k}}\right|^{2}\right)\left(1+\frac{1}{z^{2}}\right)$
[…]
where, recall that, $\bar{y}_{\mathrm{e}}=k/(a_{\mathrm{e}}H_{\mathrm{e}})$.
As before, the Bogoliubov coefficients $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$ are to be determined by matching the mode function $\mu_{k}(\eta)$ and its time derivative with those of the previous epoch at the time of transition from one epoch to another. On matching the mode function and its time derivative in the radiation and matter-dominated epochs at the time of radiation-matter equality, we find that the coefficients $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$ can be written as
$\alpha_{k}^{\mathrm{m}}$ $=\frac{1}{W_{k}^{\mathrm{m}}}\biggl\{-\left[\alpha_{k}^{\mathrm{r}}m_{k}^{\prime}(\eta_{\mathrm{eq}})+\beta_{k}^{\mathrm{r}}n_{k}^{\prime}(\eta_{\mathrm{eq}})\right]q_{k}(\eta_{\mathrm{eq}})+\left[\alpha_{k}^{\mathrm{r}}m_{k}(\eta_{\mathrm{eq}})+\beta_{k}^{\mathrm{r}}n_{k}(\eta_{\mathrm{eq}})\right]q_{k}^{\prime}(\eta_{\mathrm{eq}})\biggr\},$ (3.19a)
$\beta_{k}^{\mathrm{m}}$ $=\frac{1}{W_{k}^{\mathrm{m}}}\biggl\{\left[\alpha_{k}^{\mathrm{r}}m_{k}^{\prime}(\eta_{\mathrm{eq}})+\beta_{k}^{\mathrm{r}}n_{k}^{\prime}(\eta_{\mathrm{eq}})\right]p_{k}(\eta_{\mathrm{eq}})-\left[\alpha_{k}^{\mathrm{r}}m_{k}(\eta_{\mathrm{eq}})+\beta_{k}^{\mathrm{r}}n_{k}(\eta_{\mathrm{eq}})\right]p_{k}^{\prime}(\eta_{\mathrm{eq}})\biggr\},$ (3.19b)
where $W_{k}^{\mathrm{m}}$ is the Wronskian defined as
$W_{k}^{\mathrm{r}}(\eta)\equiv m_{k}(\eta)n_{k}^{\prime}(\eta)-m_{k}^{\prime}(\eta)n_{k}(\eta).$ (3.20)
As in the radiation-dominated epoch, the Wronskian is a constant, which can be determined to be $W_{k}^{\mathrm{m}}=4ik/\pi$. Using the forms of the functions $(m_{k},n_{k})$ and $(p_{k},q_{k})$ from Eqs. (2.16) and (3.15) in the above expressions, we obtain $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$ to be
$\alpha_{k}^{\mathrm{m}}$ $=\frac{\mathrm{e}^{ix_{\mathrm{eq}}}}{2}\left[\alpha_{k}^{\mathrm{r}}\left(-2i+\frac{1}{x_{\mathrm{eq}}}+\frac{i}{4x^{2}_{\mathrm{eq}}}\right)-\frac{i\beta_{k}^{\mathrm{r}}}{4x_{\mathrm{eq}}^{2}}\mathrm{e}^{2ix_{\mathrm{eq}}}\right],$ (3.21a)
$\beta_{k}^{\mathrm{m}}$ $=\frac{\mathrm{e}^{-ix_{\mathrm{eq}}}}{2}\left[\frac{i\alpha_{k}^{\mathrm{r}}}{4x^{2}_{\mathrm{eq}}}\mathrm{e}^{-2ix_{\mathrm{eq}}}+\beta_{k}^{\mathrm{r}}\left(2i+\frac{1}{x_{\mathrm{eq}}}-\frac{i}{4x^{2}_{\mathrm{eq}}}\right)\right],$ (3.21b)
where $x_{\mathrm{eq}}=k(\eta_{\mathrm{eq}}-\eta_{\mathrm{r}})$, and we have used the fact that $z=k(\eta+\eta_{\mathrm{eq}}-2\eta_{\mathrm{r}})$ and hence $z_{\mathrm{eq}}=2x_{\mathrm{eq}}$. The quantities involving the Bogoliubov coefficients $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$ that appear in the PS (3.16) above can be expressed as
$|\alpha_{k}^{\mathrm{m}}|^{2}$ $=|\alpha_{k}^{\mathrm{r}}|^{2}\left(1+\frac{1}{64x^{4}_{\mathrm{eq}}}\right)+\frac{|\beta_{k}^{\mathrm{r}}|^{2}}{64x^{4}_{\mathrm{eq}}}+\frac{\alpha_{k}^{\mathrm{r}}\beta_{k}^{\mathrm{r}*}}{16x^{2}_{\mathrm{eq}}}\left(2+\frac{i}{x_{\mathrm{eq}}}-\frac{1}{4x^{2}_{\mathrm{eq}}}\right)\mathrm{e}^{-2ix_{\mathrm{eq}}}$
$\quad+\frac{\alpha_{k}^{\mathrm{r}*}\beta_{k}^{\mathrm{r}}}{16x^{2}_{\mathrm{eq}}}\left(2-\frac{i}{x_{\mathrm{eq}}}-\frac{1}{4x^{2}_{\mathrm{eq}}}\right)\mathrm{e}^{2ix_{\mathrm{eq}}},$ (3.22a)
$|\beta_{k}^{\mathrm{m}}|^{2}$ $=\frac{|\alpha_{k}^{\mathrm{r}}|^{2}}{64x^{4}_{\mathrm{eq}}}+|\beta_{k}^{\mathrm{r}}|^{2}\left(1+\frac{1}{64x^{4}_{\mathrm{eq}}}\right)+\frac{\alpha_{k}^{\mathrm{r}}\beta_{k}^{\mathrm{r}}{}^{*}}{16x^{2}_{\mathrm{eq}}}\left(2+\frac{i}{x_{\mathrm{eq}}}-\frac{1}{4x^{2}_{\mathrm{eq}}}\right)\mathrm{e}^{-2ix_{\mathrm{eq}}}$
$\quad+\frac{\alpha_{k}^{\mathrm{r}}{}^{*}\beta_{k}^{\mathrm{r}}}{16x^{2}_{\mathrm{eq}}}\left(2-\frac{i}{x_{\mathrm{eq}}}-\frac{1}{4x^{2}_{\mathrm{eq}}}\right)\mathrm{e}^{2ix_{\mathrm{eq}}},$ (3.22b)
[…]
$\quad-\frac{1}{4}\left(-\frac{4}{x_{\mathrm{eq}}}+\frac{1}{2x^{3}_{\mathrm{eq}}}\right)\cos(2x_{\mathrm{eq}})-\frac{1}{64x^{4}_{\mathrm{eq}}}\sin(6x_{\mathrm{eq}})\Biggr],$ (3.22c)
[…]
$\quad-\frac{1}{4}\left(-\frac{4}{x_{\mathrm{eq}}}+\frac{1}{2x^{3}_{\mathrm{eq}}}\right)\sin(2x_{\mathrm{eq}})+\frac{1}{64x^{4}_{\mathrm{eq}}}\cos(6x_{\mathrm{eq}})\Biggr].$ (3.22d)
In the case of instantaneous transitions from de Sitter inflation to the epoch of radiation and matter domination, the PS of PGWs today can be arrived at by substituting these expressions, along with the expressions (3.4) for the Bogoliubov coefficients $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$, in Eq. (3.2). The complete expression proves to be rather long and cumbersome to write down explicitly. Therefore, let us discuss the shape of the PS in the different limits of interest.
The shape of the PS during the matter-dominated era is controlled by the three quantities (in contrast to two during the radiation-dominated era): $x_{\mathrm{e}}=y_{\mathrm{e}}$, $x_{\mathrm{eq}}=y_{\mathrm{e}}a_{\mathrm{eq}}/a_{\mathrm{e}}$ and $z_{0}=2x_{\mathrm{eq}}(a_{0}/a_{\mathrm{eq}})^{1/2}$, where, recall that, $a_{0}$ is the scale factor today. Note that these relations imply the inequality $z_{0}>x_{\mathrm{eq}}>y_{\mathrm{e}}$. Let us now consider the behavior of the PS of PGWs in the two limits $x_{\mathrm{eq}}\gg 1$ and $x_{\mathrm{eq}}\ll 1$. In the limit $x_{\mathrm{eq}}\gg 1$, it is easy to show that the quantities in the expressions (3.22) reduce to the following forms:
$|\alpha_{k}^{\mathrm{m}}|^{2}$ $\simeq|\alpha_{k}^{\mathrm{r}}|^{2},$ (3.23a)
$|\beta_{k}^{\mathrm{m}}|^{2}$ $\simeq|\beta_{k}^{\mathrm{r}}|^{2},$ (3.23b)
[…]
It then follows that, in the limit, the PS of PGWs in the matter-dominated era reduces to
$\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ $=\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\frac{1}{y_{\mathrm{e}}^{2}}\Biggl[y_{\mathrm{e}}^{4}+\frac{1}{2}-\frac{1}{2}\left(1-2y_{\mathrm{e}}^{2}\right)\cos\left(2z-2y_{\mathrm{e}}-2x_{\mathrm{eq}}\right)$
$\quad+y_{\mathrm{e}}\sin\left(2z-2y_{\mathrm{e}}-2x_{\mathrm{eq}}\right)\Biggr].$ (3.24)
Apart from the difference in the phase factors of the trigonometric functions, this PS has the same form as the PS in Eq. (3.5) we had obtained in the radiation-dominated era. From the above expression, we can easily arrive at the PS in the limits $x_{\mathrm{eq}}\gg 1$ and $y_{\rm e}\gg 1$ or, equivalently, when $k\gg k_{\mathrm{e}}$. In these limits, we obtain that
$\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)\simeq\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\left[y_{\mathrm{e}}^{2}+\cos\left(2z-2y_{\mathrm{e}}-2x_{\mathrm{eq}}\right)\right].$ (3.25)
On the other hand, in the limits $x_{\mathrm{eq}}\gg 1$ and $y_{\rm e}\ll 1$ or, equivalently, when $k_{\mathrm{eq}}\ll k\ll k_{\mathrm{e}}$, we obtain that
$\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)\simeq\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\frac{1}{y_{\mathrm{e}}^{2}}\sin^{2}(x_{\rm eq}-z).$ (3.26)
We shall now discuss the behavior of the PS over large scales such that $x_{\mathrm{eq}}\ll 1$. In the limit $x_{\mathrm{eq}}\ll 1$, from Eqs. (3.21), we obtain that
$\alpha_{k}^{\mathrm{m}}$ $\simeq\frac{i}{8x_{\mathrm{eq}}^{2}}(\alpha_{k}^{\mathrm{r}}-\beta_{k}^{\mathrm{r}})+\frac{3}{8x_{\mathrm{eq}}}(\alpha_{k}^{\mathrm{r}}+\beta_{k}^{\mathrm{r}})-\frac{9i}{16}(\alpha_{k}^{\mathrm{r}}-\beta_{k}^{\mathrm{r}})+\frac{x_{\mathrm{eq}}}{48}(37\alpha_{k}^{\mathrm{r}}-27\beta_{k}^{\mathrm{r}}),$ (3.27a)
$\beta_{k}^{\mathrm{m}}$ $\simeq\frac{i}{8x_{\mathrm{eq}}^{2}}(\alpha_{k}^{\mathrm{r}}-\beta_{k}^{\mathrm{r}})+\frac{3}{8x_{\mathrm{eq}}}(\alpha_{k}^{\mathrm{r}}+\beta_{k}^{\mathrm{r}})-\frac{9i}{16}(\alpha_{k}^{\mathrm{r}}-\beta_{k}^{\mathrm{r}})+\frac{x_{\mathrm{eq}}}{48}(-27\alpha_{k}^{\mathrm{r}}+37\beta_{k}^{\mathrm{r}}).$ (3.27b)
Since $a_{0}$ and $H_{0}$ are the scale factor and the Hubble parameter today, $k_{0}/a_{0}=H_{0}$ is the largest wave number of observational interest. On substituting the above expressions for $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$ in the PS (3.16) and considering the limit $z\gg 1$, which corresponds to $k_{0}\ll k\ll k_{\mathrm{eq}}$, we obtain that
$\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)\simeq\frac{9}{8\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\frac{1}{y_{\mathrm{e}}^{2}x_{\mathrm{eq}}^{2}}\cos^{2}(z).$ (3.28)
In contrast, if we take the limits $x_{\mathrm{eq}}\ll 1$, $y_{\mathrm{e}}\ll 1$ and $z\ll 1$ one after the another, we find that the PS (3.16) simplifies to the form
$\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)\simeq\frac{1}{8\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\frac{z^{4}}{y_{\mathrm{e}}^{2}x_{\mathrm{eq}}^{2}}=\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2},$ (3.29)
which is the standard scale-invariant behavior of the PS of PGWs on the largest scales.
We can plot the PS of PGWs today using Eq. (3.2), along with the expressions (3.21) and (3.4) for the Bogoliubov coefficients $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$ and $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$. To plot the PS, we also require the ratios $H_{\mathrm{e}}/M_{{}_{\mathrm{Pl}}}$ and $a_{\mathrm{e}}/a_{0}$. We can express the ratio $a_{\mathrm{e}}/a_{0}$ as
$\frac{a_{\mathrm{e}}}{a_{0}}=\frac{a_{\mathrm{e}}}{a_{\mathrm{eq}}}\frac{a_{\mathrm{eq}}}{a_{0}},$ (3.30)
and we have already evaluated the ratio $a_{\mathrm{e}}/a_{\mathrm{eq}}$ [cf. Eq. (3.1)]. The ratio of scale factors at the time of radiation-matter equality and today is given by
$\frac{a_{\mathrm{eq}}}{a_{0}}=\frac{\Omega_{\gamma}h^{2}+\Omega_{\nu}h^{2}}{\Omega_{\mathrm{c}}h^{2}+\Omega_{\mathrm{b}}h^{2}}.$ (3.31)
In Fig. 2, we have plotted the PS (3.2) evaluated today as a function of $y_{\mathrm{e}}$ for the same set of parameters as in the previous figure. It should be clear from the figure that the PS exhibits four different regimes, viz. a scale-invariant behavior (over $k\lesssim k_{0}$), $k^{-4}$ and $k^{-2}$ behavior with superimposed oscillations (over $k_{0}\lesssim k\lesssim k_{\mathrm{eq}}$ and $k_{\mathrm{eq}}\lesssim k\lesssim k_{\mathrm{e}}$), and a $k^{2}$ behavior (over $k\gtrsim k_{\mathrm{e}}$), as described by Eqs. (3.29), (3.28), (3.26) and (3.25), respectively.
Figure 2: The PS of PGWs $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ evaluated today, i.e. at $\eta_{0}$, has been plotted for the case of instantaneous transition from de Sitter inflation to the epoch of radiation domination and subsequently to the epoch of matter domination [cf. Eq. (3.2)]. We have plotted the PS for the three values of $g_{\ast,\mathrm{rh}}^{1/4}T_{\mathrm{rh}}$ we had plotted in the previous figure. Note that, in addition to the different behavior of the PS in the three regimes mentioned in the previous figure, there arises a fourth regime when $x_{\mathrm{eq}}\ll 1$ and $z\gg 1$, where the PS behaves as $k^{-4}$ with superimposed oscillations [cf. Eq. (3.28)]. These wave numbers correspond to $k_{0}\lesssim k\lesssim k_{\mathrm{eq}}$ which re-enter the Hubble radius during the epoch of matter domination.
3.3 Difficulties with the PS
In the previous section, we had obtained the expression for the PS of PGWs in the case of instantaneous transitions from de Sitter inflation to the epochs of radiation and matter domination. We shall now examine the expression closely and highlight some of the behavior which can be considered as physically unsatisfactory.
Figure 3: The PS of PGWs $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ evaluated today, i.e. at the conformal time $\eta_{0}$, has been plotted (in red) assuming $g_{\ast,\mathrm{rh}}^{1/4}T_{\mathrm{rh}}=1.8\times 10^{16}\,\mathrm{GeV}$ (corresponding to a tensor-to-scalar ratio of $r\simeq 0.034$, which is consistent with the current constraints) for the case of instantaneous transitions from de Sitter inflation to the epochs of radiation and matter domination. In contrast to the previous two figures, we have plotted the PS as a function of frequency $f$ rather than the dimensionless ratio $y_{\mathrm{e}}$. For the above-mentioned scenario and reheating temperature, the different wave numbers can be estimated to be $(k_{0}/a_{0},k_{\mathrm{eq}}/a_{0},k_{\mathrm{e}}/a_{0})\simeq(1.2\times 10^{-4},0.007,1.24\times 10^{23})\,\mathrm{Mpc}^{-1}$, which correspond to the frequencies $(f_{0},f_{\mathrm{eq}},f_{\mathrm{e}})=(1.94\times 10^{-19},1.14\times 10^{-17},1.92\times 10^{8})\,\mathrm{Hz}$ (indicated by the vertical dotted-dashed, dashed and dotted lines). In the figure, we have also included the sensitivity curves of various present and future GW observatories operating over a wide range of frequencies (in this context, see the discussion in the introductory section as well as Refs. [102, 76, 103]). While the curves for ADMX, SQMS, GB, JURA and IAXO represent projected future sensitivities, it is important to point out that the sensitivity curves of some detectors at high frequencies, such as BAW, ARCADE, OSQAR and CAST, are actually current upper bounds, not mere projections. The fact that CAST has not observed the signal implies that the PS of PGWs that we have plotted is ruled out. It should also be clear from the figure that, unless the $k^{2}$ rise in the PS of PGWs is truncated, the signal can, in principle, be observed by one or more of the detectors operating at high frequencies.
Previously, in Fig. 2, we had plotted the PS of PGWs today as a function of the dimensionless variable $y_{\mathrm{e}}$. Instead, in Fig. 3, we have plotted the same PS as a function of frequency ff [^3] For simplicity, we shall ignore the effects due to late-time acceleration since this affects only the largest scales or the smallest frequencies.. We have plotted the PS for the highest value of the reheating temperature $T_{\mathrm{rh}}$ that is consistent with the current bounds on the tensor-to-scalar ratio $r$ from the CMB [11, 31]. Specifically, we shall be interested in the behavior of the PS of PGWs over wave numbers such that $k\gtrsim k_{\mathrm{e}}=\eta_{\mathrm{e}}^{-1}$. Note that these wave numbers are always inside the Hubble radius, from the initial stages of inflation until today. Therefore, clearly, the PS of PGWs over this range of wave numbers reflects the unprocessed Bunch-Davies vacuum. It is then interesting to ask whether such a signal can be detected.
To answer the question, in Fig. 3, we have also included the sensitivity curves of the various ongoing and forthcoming GW observatories such as SKA, LISA, BBO, TianQin, DECIGO, CE, ET and LVK over the smaller range of frequencies (for a brief discussion on the relation between the characteristic strain in terms of which the sensitivity curves are often indicated and the sensitivity curves for the PS, see App. B). In addition, we have also included the sensitivity curves of detectors that operate (or are set to operate) at high frequencies such as BAW, ADMX, SQMS, ARCADE, GB, JURA, IAXO HET + SPD, OCSAR, JURA, IAXO and CAST. As should be evident from the figure, since the PS of PGWs rises as $k^{2}$ over $k\gtrsim k_{\mathrm{e}}$, the PS intersects the sensitivity curves of some of the detectors. In other words, it should be possible to detect the signal by one or more of the detectors at high frequencies. However, the $k^{2}$ rise in the PS on small scales is actually problematic. The reason being that, while the curves for ADMX, SQMS, GB, JURA and IAXO are projected future sensitivities, the sensitivity curves for BAW, ARCADE, OSQAR and CAST actually represent upper bounds. The fact that, contrary to what the figure suggests, no signal has been seen by CAST casts a shadow on the validity of the calculations that leads to the $k^{2}$ rise in the PS of PGWs at high frequencies. Moreover, as can be seen from Eq. (2.10), the $k^{2}$ rise would lead to a divergence in the two-point correlation function in real space. These arguments indicate that PS of PGWs needs to be regularized, a topic that we shall turn to in the next section.
4 Regularization of the PS of PGWs
In the previous section, assuming the Bunch-Davies initial conditions, we calculated the PS of PGWs at the time of radiation-matter equality as well as today. We found that, over small scales such that $y_{\mathrm{e}}\gg 1$, the PS behaves as $k^{2}$. As is well known, it is such a rise that leads to ultraviolet divergences in the coincident limit of the two-point functions of quantum fields in real space. In quantum field theory, these divergences are usually dealt with the help of regularization and renormalization procedures (for discussions in this context, see the books [104, 105, 106, 107]). For this reason, we shall also regularize the two-point function in Fourier space, i.e. the PS (for related discussions, see Refs. [84, 91, 85, 86, 87, 88, 89, 90]). We should emphasize that this last statement is non-trivial. Note that, in real space, divergences arise in the coincident limit of two-point functions of quantum fields because they involve an integral over all the wave numbers. In contrast, the two-point function in Fourier space, i.e. the PS, is evaluated at a given wave number. Therefore, actually, the PS does not contain any divergence. Nevertheless, the PS has to be regularized in order to ensure that the unphysical $k^{2}$ behavior does not arise at small scales.
4.1 Employing the method of adiabatic subtraction
The regularization procedure that is ideally suited for time-dependent situations such as the FLRW universe is the method of adiabatic regularization (for the original discussions, see Refs. [108, 109, 110, 111, 112, 113]; in this context, also see the books [106, 107]). Since we shall be interested in regularizing quantities at a given wave number, the procedure turns out to be relatively simple to employ [^4] In due course, it will be interesting to examine if other regularization procedures, such as, say, the point-splitting method, lead to the same results.. The method involves identifying the contribution to the two-point function at a given adiabatic order and subtracting it from the complete contribution. In order to parametrize the slowness of the changes in the background and calculate the corresponding contributions, it proves to be convenient to introduce a parameter, say, $T$, often referred to as the adiabatic parameter. Thereafter, an expansion in the inverse powers of $T$ is carried out and the terms at the order $T^{-n}$ is referred to as the contributions at $n$th adiabatic order [106]. Eventually, the parameter $T$ is set to unity at the end of the calculations. As we shall see, the regularization of the PS is carried out by subtracting terms up to the second adiabatic order.
Consider a field propagating in a FLRW universe. The homogeneity of the background allows us to decompose the field in terms of the Fourier modes and the isotropy of the background implies that the mode functions will depend only on the wave number $k$. In general, a suitably rescaled Fourier mode function of the field, say, $\chi_{k}$, will satisfy an equation of motion of the following form:
$\chi_{k}^{\prime\prime}(\eta)+\left[\omega_{k}^{2}(\eta)+\sigma^{2}(\eta)\right]\chi_{k}(\eta)=0,$ (4.1)
where $\omega_{k}^{2}(\eta)$ and $\sigma^{2}(\eta)$ are functions of time. For instance, in the case of a massive scalar field, the quantity $\omega_{k}^{2}(\eta)$ will depend on the wave number, the scale factor of the FLRW universe and its time derivatives, while the quantity $\sigma^{2}(\eta)$ will depend on the mass of the field and the scale factor. To solve such an equation, let us assume the WKB ansatz
[…]
where $\eta_{\mathrm{i}}$ is some initial time. On substituting this ansatz into the equation of motion (4.1), we obtain that
$W_{k}^{2}(\eta)=\omega_{k}^{2}(\eta)+\frac{\sigma^{2}(\eta)}{T^{2}}-\frac{1}{2T^{2}}\left[\frac{W_{k}^{\prime\prime}}{W_{k}}-\frac{3}{2}\left(\frac{W_{k}^{\prime}}{W_{k}}\right)^{2}\right].$ (4.3)
where we have introduced powers of $1/T$ to facilitate the adiabatic expansion. We can now expand $W_{k}(\eta)$ in adiabatic order, i.e. we can write
$W_{k}(\eta)=\sum_{n=0}^{\infty}\frac{W^{(n)}_{k}(\eta)}{T^{n}},$ (4.4)
where the index (n){(n)} refers to the adiabatic order of the expansion. In the following, we shall assume that $\omega_{k}(\eta)$ and $\sigma^{2}(\eta)$ are of adiabatic orders zero and two, respectively. On substituting this expansion into Eq. (4.3) and, on separating the terms by the adiabatic order (i.e. equating the coefficients of $T^{-n}$), we obtain that (in this context, see, for instance, Ref. [90])
$W_{k}^{(0)}$ $=\omega_{k},$ (4.5a)
$W_{k}^{(1)}$ $=0,$ (4.5b)
$W_{k}^{(2)}$ $=\frac{\sigma^{2}}{2\omega_{k}}+\frac{3\omega_{k}^{\prime 2}}{8\omega_{k}^{3}}-\frac{\omega_{k}^{\prime\prime}}{4\omega_{k}^{2}}.$ (4.5c)
For instance, note that, the adiabatic solution (4.1) leads to
$|\chi_{k}^{\mathrm{ad}}|^{2}=\frac{1}{2W_{k}}.$ (4.6)
According to Eq. (4.4), we have (when $T$ eventually set to unity)
$W_{k}^{-1}\equiv(W_{k}^{-1})^{(0)}+(W_{k}^{-1})^{(2)}+\cdots,$ (4.7)
with the terms at the different adiabatic orders $(W_{k}^{-1})^{(n)}$ defined as
$(W_{k}^{-1})^{(0)}$ $=\frac{1}{\omega_{k}},$ (4.8a)
$(W_{k}^{-1})^{(2)}$ $=-\frac{W_{k}^{(2)}}{\omega_{k}^{2}}.$ (4.8b)
4.2 Regularization of the PS of PGWs
In the case of the rescaled mode function $\mu_{k}(\eta)$ that characterizes the PGWs [cf. Eq. (2.9)], we have
$\omega_{k}$ $=k,$ (4.9a)
$\sigma^{2}(\eta)$ $=-\frac{a^{\prime\prime}}{a}.$ (4.9b)
To regularize the PS of PGWs, we need to calculate the quantity $|\mu_{k}|_{\mathrm{ad}}^{2}$ up to the second adiabatic order. We have
$W_{k}^{(0)}$ $=k,$ (4.10a)
$W_{k}^{(2)}$ $=-\frac{1}{2k}\frac{a^{\prime\prime}}{a}.$ (4.10b)
and, on using Eqs. (4.6)–(4.8), we obtain the contribution to the PS of PGWs due to the adiabatic solution up to the second order to be
$\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{ad}}(k,\eta)$ $=\frac{8}{M_{{}_{\mathrm{Pl}}}^{2}}\frac{k^{3}}{2\pi^{2}}\frac{|\mu_{k}^{\mathrm{ad}}|^{2}}{a^{2}}=\frac{8}{M_{{}_{\mathrm{Pl}}}^{2}}\frac{k^{3}}{2\pi^{2}a^{2}}\frac{1}{2W_{k}}$
$=\frac{8}{M_{{}_{\mathrm{Pl}}}^{2}}\frac{k^{3}}{2\pi^{2}a^{2}}\,\frac{1}{2W_{k}^{(0)}}\left(1-\frac{W_{k}^{(2)}}{W_{k}^{(0)}}\right)=\frac{8}{M_{{}_{\mathrm{Pl}}}^{2}}\frac{k^{3}}{2\pi^{2}a^{2}}\frac{1}{2k}\left(1+\frac{a^{\prime\prime}}{2k^{2}a}\right).$ (4.11)
We should mention that the final expression is valid for arbitrary scale factor. The above adiabatic contribution should be subtracted from the original PS to lead to the following regularized PS of PGWs:
$\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)\equiv\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)-\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{ad}}(k,\eta).$ (4.12)
Note that, in the post-inflationary universe, we can express the adiabatic contribution as follows:
$\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{ad}}(k,\eta)$ $=\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\bar{y}_{\mathrm{e}}^{2}\left(1+\frac{1}{2k^{2}}\frac{a^{\prime\prime}}{a}\right),$ (4.13)
where, recall that, $\bar{y}_{\mathrm{e}}=k/(a_{\mathrm{e}}H_{\mathrm{e}})$.
Let us now apply this procedure to the PS of PGWs we obtained during the radiation domination era. During radiation domination, we have $a^{\prime\prime}=0$. Hence, in the expression for $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{ad}}(k,\eta)$ in Eq. (4.13) above, the second term vanishes. Upon subtracting the remaining contribution from the PS of PGWs as given by Eq. (2.3), we obtain the regularized PS during the epoch of radiation domination to be
[…]
which is valid for any coefficients $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$.
Similarly, during the epoch of matter domination, the contribution to the PS of PGWs at the second adiabatic order can be written as [cf. Eqs. (4.13) and (3.12)]
$\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{ad}}(k,\eta)=\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\bar{y}_{\mathrm{e}}^{2}\left(1+\frac{1}{z^{2}}\right).$ (4.15)
On subtracting this contribution from the PS of PGWs as given by Eq. (3.2), we obtain the regularized PS during the epoch of matter domination to be
[…]
which is valid for any coefficients $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$.
4.3 Regularized PS in the case of the instantaneous transitions
In the case of the instantaneous transition from de Sitter inflation to radiation domination, the regularized PS above takes the form (with $\bar{y}_{\mathrm{e}}=y_{\mathrm{e}}$)
$\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)=\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\frac{1}{2y_{\mathrm{e}}^{2}}\biggl\{1-\left(1-2y_{\mathrm{e}}^{2}\right)\cos\left[2(x-y_{\mathrm{e}})\right]+2y_{\mathrm{e}}\sin\left[2(x-y_{\mathrm{e}})\right]\biggr\}.$ (4.17)
We find that, for wavenumbers that are in the super-Hubble domain at the end of inflation, i.e. when $y_{\mathrm{e}}\ll 1$, the above expression for $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ exactly matches the PS (3.6) on large scales that we had obtained. On the contrary, for wavenumbers that are in the sub-Hubble domain at the end of inflation, i.e. when $y_{\mathrm{e}}\gg 1$, we obtain that
$\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)\simeq\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\cos\left[2(x-y_{\mathrm{e}})\right].$ (4.18)
This should be contrasted with the result (3.8) for the unregularized PS in this limit. Clearly, there is no $k^{2}$ rise anymore for $k\gtrsim k_{\mathrm{e}}$. In fact, $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ oscillates about a vanishing mean value in this regime. In Fig. 4, we have plotted the original and the regularized PS of PGWs viz. $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ and $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ [as given by Eqs. (3.5) and (4.17)] evaluated during the early stages of the radiation-dominated epoch.
Figure 4: The actual and the regularized PS of PGWs, viz. $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ and $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$, evaluated during the early stages of the radiation-dominated epoch at $\eta=-2\eta_{\mathrm{e}}=2|\eta_{\mathrm{e}}|$ have been plotted (in red and blue) for the case of instantaneous transition from de Sitter inflation [cf. Eq. (4.17)]. We have plotted the PS for the same value of $g_{\ast,\mathrm{rh}}^{1/4}T_{\mathrm{rh}}$ as in the previous figure. Note that, over $y_{\mathrm{e}}>1$ (indicated by the vertical dotted line), while the original PS rises as $k^{2}$ (with oscillations superimposed upon it), the regularized PS oscillates about zero with a constant amplitude. In other words, over $k\gtrsim k_{\mathrm{e}}$, $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ can turn negative.
The figure highlights the points we have made above. As expected, the regularization curtails the $k^{2}$ rise that occurs in the original PS for $k\gtrsim k_{\mathrm{e}}$. However, as should be clear from the figure, the regularized PS can be negative in this regime (in this context, also see the discussion and figure in Ref. [90]). Moreover, the process of regularization does not alter the conventional (i.e. the unregularized) PS over wave numbers $k\lesssim k_{\mathrm{e}}$. The last point is reassuring since the regularized PS will not alter the standard imprints on the CMB.
Let us now turn to the regularized PS of PGWs in the matter-dominated epoch. In the case of instantaneous transitions from de Sitter inflation to radiation domination and further to matter domination, the regularized PS of PGWs can be arrived at by substituting the expressions (3.21) for the Bogoliubov coefficients $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$, along with the expressions (3.4) for $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$, in Eq. (4.2). As in the case of the original PS, the expression for the regularized PS also turns out to be rather lengthy. However, as we had seen earlier, the PS reduces to a simpler form in the limit $x_{\mathrm{eq}}\gg 1$ [cf. Eq. (3.24)]. The corresponding regularized PS can be obtained to be (since, in this limit, $z\gg 1$ as well)
$\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ $=\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\frac{1}{y_{\mathrm{e}}^{2}}\Biggl[\frac{1}{2}-\frac{1}{2}\left(1-2y_{\mathrm{e}}^{2}\right)\cos\left(2z-2y_{\mathrm{e}}-2x_{\mathrm{eq}}\right)$
$\quad+y_{\mathrm{e}}\sin\left(2z-2y_{\mathrm{e}}-2x_{\mathrm{eq}}\right)\Biggr].$ (4.19)
In the limit $y_{\mathrm{e}}\gg 1$, this regularized PS simplifies to the form
$\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)\simeq\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\cos\left(2z-2y_{\mathrm{e}}-2x_{\mathrm{eq}}\right),$ (4.20)
which, barring the phase factor, has the same behavior as in the radiation-dominated epoch [cf. Eq. (4.18)]. In Fig. 5, we have plotted the original PS and the regularized PS of PGWs evaluated today.
Figure 5: The actual and the regularized PS of PGWs, viz. $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ and $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$, evaluated today at the conformal time $\eta_{0}$, have been plotted (in red and blue) for the case of instantaneous transition from de Sitter inflation to the epochs of radiation and matter domination [cf. Eqs. (3.2) and (4.2)]. We have chosen to work with $g_{\ast,\mathrm{rh}}^{1/4}T_{\mathrm{rh}}=1.8\times 10^{16}\,\mathrm{GeV}$ as in the previous two figures. Also, as in Fig. 3, the vertical lines indicate the wave numbers $k_{0}/a_{0}$, $k_{\mathrm{eq}}/a_{0}$ and $k_{\mathrm{e}}/a_{0}$. Moreover, as in Fig. 3, we have included the sensitivity curves of the various GW observatories. It is clear that the process of regularization truncates the $k^{2}$ rise for $k\gtrsim k_{\mathrm{e}}$. Also, note that the regularization procedure does not alter the PS over scales $k\lesssim k_{\mathrm{e}}$. We find that, over $k\gtrsim k_{\mathrm{e}}$, the regularized PS $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ oscillates about zero as illustrated in the previous figure. However, the amplitude of the oscillations is significantly suppressed during the late stages of matter domination when compared to the values during the early stages of radiation domination.
4.4 Adiabatic regularization in real space
The fact that the regularized PS of PGWs $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ is negative over some scales (as we saw in the previous section) may be viewed as problematic and, in this section, we shall discuss this issue in some detail. To begin with, note that the original PS $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ is a positive definite quantity, as should be evident from Eq. (2.11). However, since the regularized PS $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ is the difference between two quantities, it does not have to be positive definite. Hence, in principle, there is no reason to expect $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ to remain positive on all scales. Secondly, it is the two-point correlation function of PGWs in real space [cf. Eq. (2.10)] that is an observable and, therefore, physical requirements should be imposed only on this quantity. The PS can be compared to the sensitivities of various GW detectors, as we have done, for instance, in Fig. 3. But, the PS of PGWs is only a derived quantity and it is inferred from the physical constraints arrived at on the two-point correlation function in real space. These constraints can imply a non-standard behavior for the regularized PS $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$. However, provided the corresponding correlation function in real space behaves in a proper fashion, the non-standard behavior of $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ need not be viewed as problematic.
Let us now highlight the reasons as to why it is the two-point correlation function in real space is the quantity that can be observationally probed. At a given spatial location, say, ${\bm{x}}=0$, the signal, say, $\mathcal{S}(t)$, observed by a GW detector can be expressed as [97, 98]
$\mathcal{S}(t)=D^{ij}h_{ij}(t),$ (4.21)
where $D^{ij}$ depends on the properties of the detector and is known as the detector tensor. Then, the mean-squared displacement of the detector can be expressed as
$\left\langle\mathcal{S}^{2}(t)\right\rangle=D^{ij}D^{mn}\left\langle h_{ij}(t)h_{mn}(t)\right\rangle,$ (4.22)
where the angular brackets denote averages over the random variable $h_{ij}$. On using the result (B.4) from App. B in the above expression, we obtain that
$\left\langle\mathcal{S}^{2}(t)\right\rangle=\frac{F}{4\pi}\int_{0}^{\infty}{\mathrm{d}}fS_{h}(f)=\frac{F}{16\pi}\int_{0}^{\infty}\frac{{\mathrm{d}}f}{f}\,\mathcal{P}_{{}_{\mathrm{T}}}(f,t),$ (4.23)
where the quantity $F$ is a number that depends on the detector, and is given by
$F=\sum_{\lambda=(+,\times)}\int_{0}^{2\pi}{\mathrm{d}}\phi\int_{0}^{\pi}{\mathrm{d}}\theta\sin\theta\left[D^{ij}\varepsilon_{ij}^{\lambda}(\hat{\bm{n}})\right]^{2}.$ (4.24)
If the GWs are of quantum origin as in the case of the PGWs of our interest, then the quantities $\mathcal{S}(t)$ and $h_{ij}$ have to be treated as quantum operators instead of random variables. In such a situation, we can define the mean-squared displacement as
$\left\langle\hat{\mathcal{S}}^{2}(\eta)\right\rangle=D^{ij}D^{mn}\left\langle\hat{h}_{ij}(\eta,{\bm{x}})\hat{h}_{mn}(\eta,{\bm{x}})\right\rangle,$ (4.25)
where the averages are to be evaluated in the quantum vacuum. Using the decomposition (2.4) of the PGWs, it is straightforward to show that, in the quantum vacuum, the mean-squared displacement can be expressed as
[…]
which is the same as the expression (4.23) in the case of stochastic and isotropic, classical GWs. It should be clear that, if the PS $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ grows as $k^{2}$ at large $k$, then the quantity $\langle\hat{\mathcal{S}}^{2}(\eta)\rangle$ is not finite, which is not physically possible. Moreover, the observable $\langle\hat{\mathcal{S}}^{2}(\eta)\rangle$ is clearly related to the two-point correlation function in real space [cf. Eq. (2.10)]. These arguments justify our claim at the beginning of this section, viz. that the observable signal is related to the correlation function in real space.
In light of the above discussion, it also becomes important to understand the implications of the procedure of adiabatic regularization in real space. The two-point correlation function of the tensor perturbations, evaluated at the same time but at different spatial locations, can be written as [cf. Eq. (2.10)]
$\delta^{im}\delta^{jn}\left\langle 0\left|\hat{h}_{mn}(\eta,{\bm{x}})\hat{h}_{ij}(\eta,{\bm{x}}^{\prime})\right|0\right\rangle=\int_{0}^{\infty}\frac{\mathrm{d}k}{k}\frac{\sin\left(k|{\bm{x}}-{\bm{x}}^{\prime}|\right)}{k|{\bm{x}}-{\bm{x}}^{\prime}|}\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta).$ (4.27)
As we have discussed before, to regularize the PS, it is the contribution at the second adiabatic order that is to be subtracted from the complete PS. In real space, the contribution to the two-point correlation function of the tensor perturbations due to the second adiabatic order in Fourier space [cf. Eq. (4.11)] can be expressed as
$\delta^{im}\delta^{jn}\left\langle 0\left|\hat{h}^{ij}(\eta,{\bm{x}})\hat{h}_{ij}(\eta,{\bm{x}}^{\prime})\right|0\right\rangle^{\mathrm{ad}}$ $=\int_{0}^{\infty}\frac{\mathrm{d}k}{k}\frac{\sin(k|{\bm{x}}-{\bm{x}}^{\prime}|)}{k|{\bm{x}}-{\bm{x}}^{\prime}|}\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{ad}}(k,\eta)$
$=\frac{2}{\pi^{2}M_{{}_{\mathrm{Pl}}}^{2}a^{2}}\left[\frac{1}{|{\bm{x}}-{\bm{x}}^{\prime}|^{2}}\int_{0}^{\infty}\mathrm{d}\kappa\,\sin\kappa+\frac{a^{\prime\prime}}{2a}\int_{0}^{\infty}\frac{{\mathrm{d}}\kappa}{\kappa^{2}}\sin\kappa\right],$ (4.28)
where we have defined $\kappa=k|{\bm{x}}-{\bm{x}}^{\prime}|$. On suitably regulating the first integral, we can evaluate it to be
$\int_{0}^{\infty}\mathrm{d}\kappa\sin\kappa=\lim_{\epsilon\to 0}\int_{0}^{\infty}\mathrm{d}\kappa\,\mathrm{e}^{-\epsilon\kappa}\sin\kappa=1.$ (4.29)
We can evaluate the second integral by introducing a comoving infrared cutoff, say, $k_{{}_{\mathrm{IR}}}$. Upon doing so, we finally arrive at the adiabatic contribution at the second order to the two-point function in real space to be
$\delta^{im}\delta^{jn}\left\langle 0\left|\hat{h}_{mn}(\eta,{\bm{x}})\hat{h}_{ij}(\eta,{\bm{x}}^{\prime})\right|0\right\rangle^{\mathrm{ad}}$ $=\frac{2}{\pi^{2}M_{{}_{\mathrm{Pl}}}^{2}a^{2}}\biggl\{\frac{1}{|{\bm{x}}-{\bm{x}}|^{2}}+\frac{a^{\prime\prime}}{2a}\biggl[\frac{\sin\left(k_{{}_{\mathrm{IR}}}|{\bm{x}}-{\bm{x}}^{\prime}|\right)}{k_{{}_{\mathrm{IR}}}|{\bm{x}}-{\bm{x}}^{\prime}|}$
$\quad-\mathrm{Ci}(k_{{}_{\mathrm{IR}}}|{\bm{x}}-{\bm{x}}^{\prime}|)\biggr]\biggr\},$ (4.30)
where $\mathrm{Ci}(z)$ denotes the cosine integral function [99]. On Taylor expanding for small $|{\bm{x}}-{\bm{x}}^{\prime}|$, we obtain that
$\delta^{im}\delta^{jn}\left\langle 0\left|\hat{h}_{mn}(\eta,{\bm{x}})\hat{h}_{ij}(\eta,{\bm{x}}^{\prime})\right|0\right\rangle^{\mathrm{ad}}$ $=\frac{2}{\pi^{2}M_{{}_{\mathrm{Pl}}}^{2}}\frac{1}{|{\bm{x}}_{\mathrm{p}}-{\bm{x}}_{\mathrm{p}}^{\prime}|^{2}}\biggl\{1+\frac{a^{\prime\prime}}{2a^{3}}|{\bm{x}}_{\mathrm{p}}-{\bm{x}}^{\prime}_{\mathrm{p}}|^{2}\biggl[1-\gamma_{{}_{\mathrm{E}}}$
$-\ln\left(k_{{}_{\mathrm{IR}}}|{\bm{x}}-{\bm{x}}^{\prime}|\right)+{\cal O}\left(k_{{}_{\mathrm{IR}}}^{2}|{\bm{x}}-{\bm{x}}^{\prime}|^{2}\right)\biggr]\biggr\},$ (4.31)
where ${\bm{x}}_{\mathrm{p}}=a{\bm{x}}$ denotes the physical distance. Note that, when $a^{\prime\prime}=0$, this quantity reduces to the correlation function of a free field in Minkowski spacetime, as expected. To regularize in real space, it is this quantity that should be subtracted from the complete correlation function.
In fact, as we already discussed, the PS should be viewed as a derived quantity, obtained from the observable quantity, which is the correlation function in real space. Indeed, the two-point correlation function of the tensor perturbations in real space can be inverted back to obtain the PS. We find that the PS of PGWs can be expressed as
$\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)=\frac{2k^{2}}{\pi}\int_{0}^{\infty}{\mathrm{d}}r\,r\sin(kr)\,\delta^{im}\delta^{jn}\left\langle 0\left|\hat{h}_{mn}(\eta,{\bm{x}})\hat{h}_{ij}(\eta,{\bm{x}}^{\prime})\right|0\right\rangle$ (4.32)
and one can easily verify that the regularized PS can similarly be written as
$\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ $=\frac{2k^{2}}{\pi}\int_{0}^{\infty}{\mathrm{d}}r\,r\sin(kr)$
$\quad\times\delta^{im}\delta^{jn}\left[\left\langle 0\left|\hat{h}_{mn}(\eta,{\bm{x}})\hat{h}_{ij}(\eta,{\bm{x}}^{\prime})\right|0\right\rangle-\left\langle 0\left|\hat{h}_{mn}(\eta,{\bm{x}})\hat{h}_{ij}(\eta,{\bm{x}}^{\prime})\right|0\right\rangle^{\mathrm{ad}}\right].$ (4.33)
From such a point of view, clearly, the fact that $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ can turn negative should not appear as problematic.
Finally, a set of last remarks are in order. Often, it has been suggested that the signal of GWs has to be averaged over time (in this context, see, for instance, Ref. [98]). We find that, if we average over time, the PS of PGWs, evaluated during the epoch of radiation domination [cf. Eq. (3.5)], reduces to the following simple form:
$\overline{\mathcal{P}}_{{}_{\mathrm{T}}}(k)=\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\frac{1}{y_{\mathrm{e}}^{2}}\left(\frac{1}{2}+y_{\mathrm{e}}^{4}\right).$ (4.34)
However, note that the averaged PS $\overline{\mathcal{P}}_{{}_{\mathrm{T}}}(k)$ behaves as $k^{-2}$ on large scales and as $k^{2}$ on small scales. In particular, it no longer leads to the scale-invariant behavior on the largest scales. Interestingly, if we average the regulated PS $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ [cf. Eq. (4.17)] during the radiation-dominated epoch, we obtain that
$\overline{\mathcal{P}}_{{}_{\mathrm{T}}}^{\,\mathrm{reg}}(k)=\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\frac{1}{y_{\mathrm{e}}^{2}},$ (4.35)
which behaves as $k^{2}$ on all scales. These suggest that averaging should be performed only on small time scales and not on cosmological time scales. It also matters whether the limits of large and small scales are considered before or after averaging. If we first take the small scale limit (i.e. the limit $k\gg k_{\mathrm{e}}$) of the regularized PS $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ [see Eq. (4.18) and (4.20)] and then carry out the averaging, then the PS vanishes identically. In other words, the resulting $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ does not turn negative on such small scales.
5 Smoothing the transition from inflation to radiation domination
It should be clear from the discussions in the previous two sections that the $k^{2}$ rise in the PS of PGWs on small scales (over $k\gtrsim k_{\mathrm{e}}$) can be circumvented by considering the regularized version of the quantity. However, when transitions are involved, the process of regularization does not prove to be sufficient to ensure that generic two-point functions in real space are well behaved. Note that, we have considered the transition from inflation to the post-inflationary, radiation-dominated epoch (and, later, to the matter-dominated epoch) to be instantaneous or abrupt. Such instantaneous transitions can be considered to be unphysical. Realistic transitions are expected to be smooth (to be precise, infinitely continuous) and, clearly, it would be interesting to investigate the effects of such smooth transitions on the PS of PGWs (for a recent discussion in this context, see Ref. [95]). In this section, we shall consider a smoother transition from inflation to the epoch of radiation domination and examine the corresponding behavior of the PS of PGWs.
Instead of considering an instantaneous transition from inflation to the epoch of radiation domination, let us introduce a smoother transition which interpolates between these two phases of evolution of the universe. As we discussed, during inflation and the post-inflationary phase, the scale factors are given by Eqs. (2.18) and (2.12). Soon, we will specialize our consideration to the case of $w=1/3$ during the post-inflationary phase, which corresponds to the radiation-dominated universe. But, for the moment, let us present equations that are valid for any value of the equation-of-state parameter $w$. Recall that inflation ends at $\eta_{\mathrm{e}}$. Let us assume that the transition occurs over the time period $\Delta\eta$. If we interpolate between the inflation and the post-inflationary phase using a linear function for $U(\eta)=a^{\prime\prime}/a$, we obtain that
$U(\eta)=\left\{\begin{array}[]{ll}\!\frac{q\,(q-1)}{\eta^{2}},&\mbox{for}\penalty 10000\ \eta\leq\eta_{\mathrm{e}},\\ \!\frac{2}{\eta_{\mathrm{e}}^{2}\,\Delta\eta}\!\left\{\left[\frac{\gamma(w)}{(1-\Delta\eta/|\eta_{\mathrm{e}}|+\eta_{w}/|\eta_{\mathrm{e}}|)^{2}}-\frac{q\,(q-1)}{2}\right](\eta-\eta_{\mathrm{e}})+\frac{q\,(q-1)}{2}\,\Delta\eta\right\},&\mbox{for}\penalty 10000\ \eta_{\mathrm{e}}\leq\eta\leq\eta_{\mathrm{e}}+\Delta\eta,\\ \!\frac{2\,\gamma(w)}{(\eta-\eta_{w})^{2}},&\mbox{for}\penalty 10000\ \eta\geq\eta_{\mathrm{e}}+\Delta\eta.\end{array}\right$. (5.4)
It is straightforward to check that the quantity $U(\eta)$ is continuous at both $\eta_{\mathrm{e}}$ and $\eta_{\mathrm{e}}+\Delta\eta$. We have plotted the $U(\eta)$ above for the case of transition from de Sitter inflation to the epoch of radiation domination in Fig. 6.
Figure 6: The behavior of the quantity $U(\eta)=a^{\prime\prime}/a$ has been plotted in the case of transition from de Sitter inflation (corresponding to q=1q=-1) to the epoch of radiation domination (corresponding to $w=1/3$), when $U(\eta)$ has been smoothed with a linear function [cf. Eq. (5.4)]. Note that we have set $\Delta\eta=2|\eta_{\mathrm{e}}|$ in plotting the figure. We have also indicated the beginning and the end of the smooth transition (as vertical dotted and dashed lines, respectively).
Evidently, while we know the forms of the scale factor during inflation and the phase dominated by the fluid or field with the equation-of-state parameter $w$, we do not know it during the intermediate transitory phase. However, since $U(\eta)=a^{\prime\prime}/a$, postulating a $U(\eta)$ actually corresponds to specifying a scale factor. Let us determine the scale factor during the transition. The differential equation satisfied by the scale factor is given by
$a^{\prime\prime}-U(\eta)a=0.$ (5.5)
which, for the above choice of $U(\eta)$, assumes the form
$a^{\prime\prime}-\frac{2}{\eta_{\mathrm{e}}^{2}\,\Delta\eta}\left\{\left[\frac{\gamma(w)}{(1-\Delta\eta/|\eta_{\mathrm{e}}|+\eta_{w}/|\eta_{\mathrm{e}}|)^{2}}-\frac{q\,(q-1)}{2}\right](\eta-\eta_{\mathrm{e}})+\frac{q\,(q-1)}{2}\,\Delta\eta\right\}\,a=0.$ (5.6)
If we define $\rho(\eta)=\alpha\,U(\eta)$, where $\alpha$ is a constant to be determined, then the above differential equation for $a(\eta)$ can be written as
$\alpha^{3}\left(\frac{\mathrm{d}U}{\mathrm{d}\eta}\right)^{2}\frac{\mathrm{d}^{2}a}{\mathrm{d}\rho^{2}}-\rho a=0.$ (5.7)
Therefore, if we choose $\alpha$ such that $\alpha^{3}({\mathrm{d}}U/{\mathrm{d}}\eta)^{2}=1$, i.e. $\alpha=\left|\mathrm{d}U/\mathrm{d}\eta\right|^{-2/3}$, then we find that the scale factor obeys the Airy differential equation. Notice that
$\left|\frac{\mathrm{d}U}{\mathrm{d}\eta}\right|=\frac{1}{|\eta_{\mathrm{e}}|^{3}}X(\eta_{\mathrm{e}},\Delta\eta,\eta_{w}),$ (5.8)
where we have defined the quantity $X$ to be
$X(\eta_{\mathrm{e}},\Delta\eta,\eta_{w})\equiv\left(\frac{2|\eta_{\mathrm{e}}|}{\Delta\eta}{}\right)\left|\frac{\gamma(w)}{(1-\Delta\eta/|\eta_{\mathrm{e}}|+\eta_{w}/|\eta_{\mathrm{e}}|)^{2}}-\frac{q(q-1)}{2}\right|$ (5.9)
which, a priori, is a factor of order unity. As a result, the quantity $\alpha$ can be written as
$\alpha=|\eta_{\mathrm{e}}|^{2}X^{-2/3}.$ (5.10)
It follows from the above considerations that the solution for the scale factor can be written in terms of Airy functions $\mathrm{Ai}(x)$ and $\mathrm{Bi}(x)$ as follows:
$\frac{a(\eta)}{a_{\mathrm{e}}}$ $=A_{1}\,u(\eta)+A_{2}\,v(\eta),$ (5.11)
where $u(\eta)=\mathrm{Ai}\left[\rho(\eta)\right]$, $v(\eta)=\mathrm{Bi}\left[\rho(\eta)\right]$ and, recall that, $a_{\mathrm{e}}=a(\eta_{\mathrm{e}})$. It is useful to note that the above solution also implies that the conformal Hubble parameter $\mathcal{H}$ during the transition is given by
${\cal H}(\eta)=\frac{\mathrm{d}\rho}{\mathrm{d}\eta}\frac{A_{1}(\mathrm{d}u/\mathrm{d}\rho)+A_{2}(\mathrm{d}v/\mathrm{d}\rho)}{A_{1}u+A_{2}v}.$ (5.12)
The two constants $A_{1}$ and $A_{2}$ are to be determined by demanding that the scale factor and its first time derivative are continuous at the end of inflation, i.e. at $\eta_{\mathrm{e}}$. These conditions lead to
$A_{1}$ $=\pi\left[\frac{\mathrm{d}v}{\mathrm{d}\rho}\biggl|_{\eta_{\mathrm{e}}}-\left(\frac{\mathrm{d}\rho}{\mathrm{d}\eta}\right)^{-1}{\cal H}_{\mathrm{e}}v(\eta_{\mathrm{e}})\right],$ (5.13a)
$A_{2}$ $=-\pi\left[\frac{\mathrm{d}u}{\mathrm{d}\rho}\biggl|_{\eta_{\mathrm{e}}}-\left(\frac{\mathrm{d}\rho}{\mathrm{d}\eta}\right)^{-1}{\cal H}_{\mathrm{e}}u(\eta_{\mathrm{e}})\right],$ (5.13b)
where $\mathcal{H}_{\mathrm{e}}=\mathcal{H}(\eta_{\mathrm{e}})$. We should mention that, in arriving at these expressions for $A_{1}$ and $A_{2}$, we have made use of the following Wronskian condition [99]:
$\mathrm{Ai}(x)\frac{{\mathrm{d}}\mathrm{Bi}(x)}{{\mathrm{d}}x}-\frac{{\mathrm{d}}\mathrm{Ai}(x)}{{\mathrm{d}}x}\mathrm{Bi}(x)=\frac{1}{\pi}.$ (5.14)
Note that the factor $\mathrm{d}\rho/\mathrm{d}\eta$ can be expressed as
$\frac{\mathrm{d}\rho}{\mathrm{d}\eta}=\biggl|\frac{\mathrm{d}U}{\mathrm{d}\eta}\biggr|^{-2/3}\frac{\mathrm{d}U}{\mathrm{d}\eta}=\varepsilon\biggl|\frac{\mathrm{d}U}{\mathrm{d}\eta}\biggr|^{1/3}=\varepsilon|\eta_{\mathrm{e}}|^{-1}X^{1/3},$ (5.15)
where $\varepsilon$ is given by
$\varepsilon=\mathrm{sgn}\left[\frac{\gamma(w)}{(1-\Delta\eta/|\eta_{\mathrm{e}}|+\eta_{w}/|\eta_{\mathrm{e}}|)^{2}}-\frac{q\,(q-1)}{2}\right].$ (5.16)
In fact, $\varepsilon$ depends on whether $\mathrm{d}U/\mathrm{d}\eta>0$ ($\varepsilon=1$) or $\mathrm{d}U/\mathrm{d}\eta<0$ ($\varepsilon=-1$). Indeed, the sign of $\varepsilon$ in the space $(q,w,\eta_{\mathrm{e}},\Delta\eta)$ is rather non-trivial. Moreover, the coefficients $A_{1}$ and $A_{2}$ involve the functions $u(\eta)$ and $v(\eta)$ evaluated at the end of inflation which, in turn, requires the calculation of $\rho(\eta_{\mathrm{e}})$. Using the relation, $\rho(\eta_{\mathrm{e}})=\alpha U(\eta_{\mathrm{e}})=q(q-1)X^{-2/3}$, we can see that $\rho(\eta_{\mathrm{e}})$ is, a priori, a quantity of order unity. Therefore, we cannot use the asymptotic behavior of the Airy functions. Explicitly, the coefficients $A_{1}$ and $A_{2}$ can be expressed as
$A_{1}$ $=\pi\left\{\mathrm{Bi}^{\prime}\left[q(q-1)X^{-2/3}\right]+\varepsilon qX^{-1/3}\,\mathrm{Bi}\left[q(q-1)X^{-2/3}\right]\right\},$ (5.17a)
$A_{2}$ $=-\pi\left\{\mathrm{Ai}^{\prime}\left[q(q-1)X^{-2/3}\right]+\varepsilon qX^{-1/3}\,\mathrm{Ai}\left[q(q-1)X^{-2/3}\right]\right\}.$ (5.17b)
where $\mathrm{Ai}^{\prime}(x)={\mathrm{d}}\mathrm{Ai}(x)/{\mathrm{d}}x$ and $\mathrm{Bi}(x)={\mathrm{d}}\mathrm{Bi}(x)/{\mathrm{d}}x$, i.e. the overprimes on the Airy functions in these two equations denote differentiation with respect to the arguments.
The discussion above was valid for any value of $w$. However, in this work, we are mainly interested in smoothing the transition from inflation to the epoch of radiation domination. For this reason, from now on, we shall set $w=1/3$. In such a case $\gamma(w)=0$ and $\mathrm{d}U/\mathrm{d}\eta<0$, and we have
$\varepsilon=\mathrm{sgn}\left[-\frac{q\,(q-1)}{2}\right]=-1.$ (5.18)
Also, we have
$X=\frac{q\,(q-1)\,|\eta_{\mathrm{e}}|}{\Delta\eta}$ (5.19)
and, note that, $X$ no longer depends on $\eta_{w}$. As a consequence, the coefficients $A_{1}$ and $A_{2}$ take the form
$\frac{A_{1}}{\pi}$ $=\mathrm{Bi}^{\prime}\left\{\left[q(q-1)\right]^{1/3}\left(\frac{\Delta\eta}{|\eta_{\mathrm{e}}|}\right)^{2/3}\right\}$
$\quad-q\left[q(q-1)\right]^{-1/3}\left(\frac{\Delta\eta}{|\eta_{\mathrm{e}}|}\right)^{1/3}\mathrm{Bi}\left\{\left[q(q-1)\right]^{1/3}\left(\frac{\Delta\eta}{|\eta_{\mathrm{e}}|}\right)^{2/3}\right\},$ (5.20a)
$\frac{A_{2}}{\pi}$ $=-\mathrm{Ai}^{\prime}\left\{\left[q(q-1)\right]^{1/3}\left(\frac{\Delta\eta}{|\eta_{\mathrm{e}}|}\right)^{2/3}\right\}$
$\quad+q\left[q(q-1)\right]^{-1/3}\left(\frac{\Delta\eta}{|\eta_{\mathrm{e}}|}\right)^{1/3}\mathrm{Ai}\left\{\left[q(q-1)\right]^{1/3}\left(\frac{\Delta\eta}{|\eta_{\mathrm{e}}|}\right)^{2/3}\right\},$ (5.20b)
and we should stress again that the overprimes on the Airy functions in these two equations denote differentiation with respect to the arguments.
We must now match the scale factor and the conformal Hubble parameter at the time $\eta_{\mathrm{e}}+\Delta\eta$, i.e. at the end of the smoother transition and the beginning of the epoch of radiation domination. These matching conditions lead to the determination of the constants $a_{\mathrm{r}}$ and $\eta_{\mathrm{r}}$ appearing in the expression (3.1b) for the scale factor during the epoch of radiation domination. When $\gamma(w)=0$, $\rho(\eta_{\mathrm{e}}+\Delta\eta)=0$ and the results [99]
$\mathrm{Ai}(0)$ $=\frac{1}{3^{2/3}\,\Gamma(2/3)},$ (5.21a)
$\mathrm{Bi}(0)$ $=\frac{1}{3^{1/6}\,\Gamma(2/3)},$ (5.21b)
$\mathrm{Ai}^{\prime}(0)$ $=-\frac{1}{3^{1/3}\,\Gamma(1/3)},$ (5.21c)
$\mathrm{Bi}^{\prime}(0)$ $=\frac{3^{1/6}}{\Gamma(1/3)},$ (5.21d)
where $\Gamma(x)$ is the Gamma function, can be used to arrive at the following expressions for $\eta_{\mathrm{r}}$ and $a_{\mathrm{r}}$:
$\frac{\eta_{\mathrm{r}}}{|\eta_{\mathrm{e}}|}$ $=\frac{\Delta\eta}{|\eta_{\mathrm{e}}|}-1-\left[q(q-1)\right]^{-1/3}\left(\frac{\Delta\eta}{|\eta_{\mathrm{e}}|}\right)^{1/3}\frac{\Gamma(1/3)}{3^{1/3}\Gamma(2/3)}\left(\frac{A_{1}+\sqrt{3}A_{2}}{A_{1}-\sqrt{3}A_{2}}\right),$ (5.22a)
$a_{\mathrm{r}}$ $=\frac{\left[q(q-1)\right]^{1/3}}{3^{1/3}\Gamma(1/3)|\eta_{\mathrm{e}}|}\left(\frac{\Delta\eta}{|\eta_{\mathrm{e}}|}\right)^{-1/3}\left(A_{1}-\sqrt{3}A_{2}\right)a_{\mathrm{e}}.$ (5.22b)
These two quantities depend only on the three parameters: $q$, $\eta_{\mathrm{e}}$ and $\Delta\eta/|\eta_{\mathrm{e}}|$.
It is also interesting to examine the manner in which the case of the instantaneous transition is recovered from the above expressions for $a_{\mathrm{r}}$ and $\eta_{\mathrm{r}}$. Clearly, the instantaneous transition corresponds to the limit $\Delta\eta\rightarrow 0$, while $\eta_{\mathrm{e}}$ remains finite. In such a limit, we find that the constants $A_{1}$ and $A_{2}$ simplify to
$A_{1}$ $\simeq\frac{3^{1/6}\pi}{\Gamma(1/3)}-\frac{\pi q\Delta\eta^{1/3}}{3^{1/6}\Gamma(2/3)[q(q-1)]^{1/3}|\eta_{\mathrm{e}}|^{1/3}},$ (5.23a)
$\sqrt{3}\,A_{2}$ $\simeq\frac{3^{1/6}\,\pi}{\Gamma(1/3)}+\frac{\pi q\Delta\eta^{1/3}}{3^{1/6}\Gamma(2/3)[q(q-1)]^{1/3}|\eta_{\mathrm{e}}|^{1/3}},$ (5.23b)
and, hence,
$A_{1}+\sqrt{3}A_{2}$ $\simeq\frac{2\,3^{1/6}\pi}{\Gamma(1/3)},$ (5.24a)
$A_{1}-\sqrt{3}A_{2}$ $\simeq-\frac{2\pi q\Delta\eta^{1/3}}{3^{1/6}\Gamma(2/3)[q(q-1)]^{1/3}|\eta_{\mathrm{e}}|^{1/3}}.$ (5.24b)
Since $\Gamma(1/3)\Gamma(2/3)=2\pi/\sqrt{3}$, the above expressions immediately lead to $a_{\mathrm{r}}=qa_{\mathrm{e}}/\eta_{\mathrm{e}}$ and $\eta_{\mathrm{r}}=(q-1)\eta_{\mathrm{e}}/q$, as we had seen earlier [see our discussion following Eq. (3.1b)].
6 Evolution of PGWs across the smoother transition
Having obtained the solution for the background across the smooth linear transition, let us now construct the solution for the tensor mode functions.
6.1 Evolution during the smoother transition
We are now in a position to solve the equation (2.9) for the rescaled mode function $\mu_{k}(\eta)$. The solutions during inflation and the radiation dominated era are already known and, hence, let us focus on the solution during the transition. If we define
$\tau(\eta)=-\alpha\left[k^{2}-U(\eta)\right]=-\alpha k^{2}+\rho(\eta),$ (6.1)
then the equation governing the rescaled mode function during the transition takes the form
$\frac{\mathrm{d}^{2}\mu_{k}}{\mathrm{d}\tau^{2}}-\tau\mu_{k}=0,$ (6.2)
which can be recognized again as the Airy differential equation. As a result, the general solution for the rescaled mode function can be expressed as
$\mu_{k}(\eta)=B_{1}i_{k}(\eta)+B_{2}j_{k}(\eta)=B_{1}\mathrm{Ai}[\tau(\eta)]+B_{2}\mathrm{Bi}[\tau(\eta)],$ (6.3)
where $\mathrm{Ai}(x)$ and $\mathrm{Bi}(x)$ are the Airy functions. The two constants $B_{1}$ and $B_{2}$ are to be determined by requiring the continuity of the rescaled mode function $\mu_{k}(\eta)$ and its time derivative at the end of inflation, i.e. at $\eta_{\mathrm{e}}$. The constants can be expressed as
$B_{1}$ $=\frac{A_{k}}{W_{k}^{\mathrm{slt}}}\left[f_{k}(\eta_{\mathrm{e}})\frac{\mathrm{d}j_{k}}{\mathrm{d}\eta}\biggl|_{\eta_{\mathrm{e}}}-\frac{\mathrm{d}f_{k}}{\mathrm{d}\eta}\biggl|_{\eta_{\mathrm{e}}}j_{k}(\eta_{\mathrm{e}})\right]$
$=\frac{A_{k}}{W_{k}^{\mathrm{slt}}}\frac{\mathrm{d}\rho}{\mathrm{d}\eta}\left[f_{k}(\eta_{\mathrm{e}})\frac{\mathrm{d}j_{k}}{\mathrm{d}\tau}\biggl|_{\eta_{\mathrm{e}}}+k\frac{\mathrm{d}f_{k}}{\mathrm{d}y}\biggl|_{\eta_{\mathrm{e}}}\left(\frac{\mathrm{d}\rho}{\mathrm{d}\eta}\right)^{-1}j_{k}(\eta_{\mathrm{e}})\right],$ (6.4a)
$B_{2}$ $=-\frac{A_{k}}{W_{k}^{\mathrm{slt}}}\left[f_{k}(\eta_{\mathrm{e}})\frac{\mathrm{d}i_{k}}{\mathrm{d}\eta}\biggl|_{\eta_{\mathrm{e}}}-\frac{\mathrm{d}f_{k}}{\mathrm{d}\eta}\biggl|_{\eta_{\mathrm{e}}}i_{k}(\eta_{\mathrm{e}})\right]$
$=-\frac{A_{k}}{W_{k}^{\mathrm{slt}}}\frac{\mathrm{d}\rho}{\mathrm{d}\eta}\left[f_{k}(\eta_{\mathrm{e}})\frac{\mathrm{d}i_{k}}{\mathrm{d}\tau}\biggl|_{\eta_{\mathrm{e}}}+k\frac{\mathrm{d}f_{k}}{\mathrm{d}y}\biggl|_{\eta_{\mathrm{e}}}\left(\frac{\mathrm{d}\rho}{\mathrm{d}\eta}\right)^{-1}i_{k}(\eta_{\mathrm{e}})\right],$ (6.4b)
where $f_{k}(\eta)$ is the mode function during inflation [cf. Eqs. (2.19) and (2.20)], and we have made use of the fact that $\mathrm{d}\tau/\mathrm{d}\eta=\mathrm{d}\rho/\mathrm{d}\eta$, since $\tau(\eta)$ and $\rho(\eta)$ differ only by a constant. The quantity $W_{k}^{\mathrm{slt}}$ is the Wronskian given by
$W_{k}^{\mathrm{slt}}=i_{k}\frac{{\mathrm{d}}j_{k}}{{\mathrm{d}}\eta}-j_{k}\frac{\mathrm{d}i_{k}}{\mathrm{d}\eta}=\frac{1}{\pi}\frac{\mathrm{d}\tau}{\mathrm{d}\eta}=\frac{\alpha}{\pi}\frac{\mathrm{d}U}{\mathrm{d}\eta}=\frac{1}{\pi}\left|\frac{\mathrm{d}U}{\mathrm{d}\eta}\right|^{-2/3}\frac{\mathrm{d}U}{\mathrm{d}\eta}=-\frac{1}{\pi}\left|\frac{\mathrm{d}U}{\mathrm{d}\eta}\right|^{1/3},$ (6.5)
where we have made use of the Wronskian condition (5.14) for the Airy functions.
To proceed further and establish the detailed form of the coefficients $B_{1}$ and $B_{2}$, we need the exact forms of the mode functions $f_{k}(\eta)$, $i_{k}(\eta)$ and $j_{k}(\eta)$. Recall that the rescaled mode function during inflation can be expressed in terms of the Hankel functions [cf. Eqs. (2.19) and (2.20)]. It proves to be convenient to rewrite the mode function $f_{k}(\eta)$ as
$f_{k}(\eta)=\sqrt{\frac{2}{\pi}}\mathrm{e}^{i\left[y-\pi(1-q)/2\right]}T_{1/2-q}(y),$ (6.6)
where the function $T_{\mu}(y)$ is defined by the following expression:
$T_{\mu}(y)=\sum_{m=0}^{\infty}\frac{(-1)^{m}}{m!}\frac{\Gamma(1/2+\mu+m)}{\Gamma(1/2+\mu-m)}\frac{1}{(2iy)^{m}},$ (6.7)
and we recall that $y=-k\eta>0$. As a consequence, we can also write that
$\frac{\mathrm{d}f_{k}}{\mathrm{d}\eta}=-k\frac{\mathrm{d}f_{k}}{\mathrm{d}y}=-k\left[i+T^{\prime}_{1/2-q}(y)/T_{1/2-q}(y)\right]f_{k},$ (6.8)
where $T^{\prime}_{1/2-q}(y)$ denotes the derivative of the function $T_{1/2-q}(y)$ with respect to its argument. Note that the mode functions $i_{k}(\eta)$ and $j_{k}(\eta)$ are described by the Airy functions with the argument $\tau(\eta)$ [see Eq. (6.3)]. The quantity $\tau_{\mathrm{e}}$ can be expressed as
$\tau_{\mathrm{e}}=-\alpha\left[k^{2}-U(\eta_{\mathrm{e}})\right]=-\alpha k^{2}+\rho(\eta_{\mathrm{e}})=-y_{\mathrm{e}}^{2}X^{-2/3}\left[1-\frac{q(q-1)}{y_{\mathrm{e}}^{2}}\right]$ (6.9)
and, in the limit $y_{\mathrm{e}}\gg 1$, $\tau(\eta_{\mathrm{e}})$ is a large, negative quantity. As a consequence, it is convenient to use the following exact expressions for the Airy functions [114]:
$\mathrm{Ai}(s)$ $=\frac{(-s)^{-1/4}}{\sqrt{\pi}}\left[A(s)\sin\left(\zeta+\frac{\pi}{4}\right)-B(s)\cos\left(\zeta+\frac{\pi}{4}\right)\right],$ (6.10a)
$\mathrm{Ai}^{\prime}(s)$ $=\frac{(-s)^{1/4}}{\sqrt{\pi}}\left[-C(s)\cos\left(\zeta+\frac{\pi}{4}\right)-D(s)\sin\left(\zeta+\frac{\pi}{4}\right)\right],$ (6.10b)
$\mathrm{Bi}(s)$ $=\frac{(-s)^{-1/4}}{\sqrt{\pi}}\left[A(s)\cos\left(\zeta+\frac{\pi}{4}\right)+B(s)\sin\left(\zeta+\frac{\pi}{4}\right)\right],$ (6.10c)
$\mathrm{Bi}^{\prime}(s)$ $=\frac{(-s)^{1/4}}{\sqrt{\pi}}\left[C(s)\sin\left(\zeta+\frac{\pi}{4}\right)-D(s)\cos\left(\zeta+\frac{\pi}{4}\right)\right],$ (6.10d)
where $\zeta=(2/3)(-s)^{3/2}$, which is valid for negative arguments, i.e. when $s<0$. The four functions $A(s)$, $B(s)$, $C(s)$ and $D(s)$ are given by [114]
$A(s)$ $=\sum_{m=0}^{\infty}(-1)^{m}\frac{c_{2m}}{\zeta^{2m}},$ (6.11a)
$B(s)$ $=\sum_{m=0}^{\infty}(-1)^{m}\frac{c_{2m+1}}{\zeta^{2m+1}},$ (6.11b)
$C(s)$ $=\sum_{m=0}^{\infty}(-1)^{m}\frac{d_{2m}}{\zeta^{2m}},$ (6.11c)
$D(s)$ $=\sum_{m=0}^{\infty}(-1)^{m}\,\frac{d_{2m+1}}{\zeta^{2m+1}},$ (6.11d)
where the coefficients $c_{m}$ and $d_{m}$ are given by
$c_{m}$ $=\frac{\Gamma(3m+1/2)}{54^{m}m!\,\Gamma(m+1/2)},$ (6.12a)
$d_{m}$ $=-\frac{6m+1}{6m-1}c_{m}.$ (6.12b)
On substituting these formulae into the expression for $B_{1}$ derived above, we obtain the following, exact result:
$B_{1}$ $=\frac{A_{k}}{W_{\mathrm{slt}}}\frac{\mathrm{d}\rho}{\mathrm{d}\eta}\left\{f_{k}(\eta_{\mathrm{e}})\frac{\mathrm{d}j_{k}}{\mathrm{d}\tau}\biggl|_{\eta_{\mathrm{e}}}+k\left[i+\frac{T^{\prime}_{1/2-q}(y_{\mathrm{e}})}{T_{1/2-q}(y_{\mathrm{e}})}\right]f_{k}(\eta_{\mathrm{e}})\left(\frac{\mathrm{d}\rho}{\mathrm{d}\eta}\right)^{-1}j_{k}(\eta_{\mathrm{e}})\right\}$
$=\frac{A_{k}}{W_{\mathrm{slt}}}\frac{\mathrm{d}\rho}{\mathrm{d}\eta}\frac{1}{\sqrt{\pi}}f_{k}(\eta_{\mathrm{e}})(-\tau_{\mathrm{e}})^{1/4}\Biggl\{C(\tau_{\mathrm{e}})\sin\left(\zeta_{\mathrm{e}}+\frac{\pi}{4}\right)-D(\tau_{\mathrm{e}})\cos\left(\zeta_{\mathrm{e}}+\frac{\pi}{4}\right)$
$\quad+k\left[i+\frac{T^{\prime}_{1/2-q}(y_{\mathrm{e}})}{T_{1/2-q}(y_{\mathrm{e}})}\right]\left(\frac{\mathrm{d}\rho}{\mathrm{d}\eta}\right)^{-1}(-\tau_{\mathrm{e}})^{-1/2}\left[A(\tau_{\mathrm{e}})\cos\left(\zeta_{\mathrm{e}}+\frac{\pi}{4}\right)+B(\tau_{\mathrm{e}})\sin\left(\zeta_{\mathrm{e}}+\frac{\pi}{4}\right)\right]\Biggr\},$ (6.13)
where we have used the expressions for the Airy functions given above. Then, if we use Eq. (5.15) and the expression for $\tau_{\mathrm{e}}$ [see Eq. (6.9)], it is easy to establish that
$\left(\frac{\mathrm{d}\rho}{\mathrm{d}\eta}\right)^{-1}(-\tau_{\mathrm{e}})^{-1/2}=\frac{\varepsilon}{k}\left[1-\frac{q(q-1)}{y_{\mathrm{e}}^{2}}\right]^{-1/2},$ (6.14)
where $\varepsilon=\pm 1$ depending on the value of the equation-of-state parameter $w$ in the subsequent epoch [cf. Eq. (5.16)]. (Recall that $\varepsilon=-1$ in the case of transition to a radiation dominated epoch.) Finally, on using the expressions (6.5) for the Wronskian and the inflationary mode function (6.6), we find that the coefficient $B_{1}$ can be written as
$B_{1}$ $=A_{k}\sqrt{2}\,\mathrm{e}^{i\left[y_{\mathrm{e}}-\pi(1-q)/2\right]}y_{\mathrm{e}}^{1/2}X^{-1/6}\left[{\cal P}_{\mathrm{s}1}\sin\left(\zeta_{\mathrm{e}}+\frac{\pi}{4}\right)+{\cal P}_{\mathrm{c}1}\cos\left(\zeta_{\mathrm{e}}+\frac{\pi}{4}\right)\right],$ (6.15)
where the quantities $({\cal P}_{\mathrm{s}1},{\cal P}_{\mathrm{c}1})$ are given by
${\cal P}_{\mathrm{s}1}$ $=\varepsilon\left[iT_{1/2-q}(y_{\mathrm{e}})+T^{\prime}_{1/2-q}(y_{\mathrm{e}})\right]\left[1-\frac{q(q-1)}{y_{\mathrm{e}}^{2}}\right]^{-1/4}B(\tau_{\mathrm{e}})$
$\quad+T_{1/2-q}(y_{\mathrm{e}})\left[1-\frac{q(q-1)}{y_{\mathrm{e}}^{2}}\right]^{1/4}C(\tau_{\mathrm{e}}),$ (6.16a)
${\cal P}_{\mathrm{c}1}$ $=\varepsilon\left[iT_{1/2-q}(y_{\mathrm{e}})+T^{\prime}_{1/2-q}(y_{\mathrm{e}})\right]\left[1-\frac{q(q-1)}{y_{\mathrm{e}}^{2}}\right]^{-1/4}A(\tau_{\mathrm{e}})$
$\quad-T_{1/2-q}(y_{\mathrm{e}})\left[1-\frac{q(q-1)}{y_{\mathrm{e}}^{2}}\right]^{1/4}D(\tau_{\mathrm{e}}),$ (6.16b)
We should stress that that the result (6.15) for $B_{1}$ is exact. Since
$\zeta_{\mathrm{e}}=\frac{2y_{\mathrm{e}}^{3}}{3X}\left[1-\frac{q(q-1)}{y_{\mathrm{e}}^{2}}\right]^{3/2},$ (6.17)
the quantity $B_{1}/A_{k}$ is in fact a function only of $y_{\mathrm{e}}$. Moreover, note that the coefficients ${\cal P}_{\mathrm{c}1}$ and ${\cal P}_{\mathrm{s}1}$, which multiply the cosine and sine functions, involve sums of negative powers of $y_{\mathrm{e}}$. Hence, they are easy to calculate at any order in the inverse powers of $y_{\mathrm{e}}$, which is exactly what is needed in order to evaluate $B_{1}$ in the small scale limit. We find that, in the limit $y_{\mathrm{e}}\gg 1$, the quantities $({\cal P}_{\mathrm{s}1},{\cal P}_{\mathrm{c}1})$ can be expressed as
${\cal P}_{\mathrm{s}1}$ $=1+\frac{iq(q-1)}{2y_{\mathrm{e}}}-\frac{q^{2}(q-1)^{2}}{8y_{\mathrm{e}}^{2}}$
$\quad+\frac{i}{48y_{\mathrm{e}}^{3}}\left[12q-10q^{2}-q^{3}(3+q-3q^{2}-q^{3})+5\varepsilon X\right]+{\cal O}\left(\frac{1}{y_{\mathrm{e}}^{4}}\right),$ (6.18a)
${\cal P}_{\mathrm{c}1}$ $=i\varepsilon-\frac{\varepsilon q(q-1)}{2y_{\mathrm{e}}}-\frac{i\varepsilon q^{2}(q-1)^{2}}{8y_{\mathrm{e}}^{2}}$
$\quad-\frac{1}{48y_{\mathrm{e}}^{3}}\left\{\left[12q-14q^{2}+q^{3}(3+q-3q^{2}+q^{3})\right]\varepsilon+7X\right\}+{\cal O}\left(\frac{1}{y_{\mathrm{e}}^{4}}\right).$ (6.18b)
In these expressions, we have expanded the quantities $({\cal P}_{\mathrm{s}1},{\cal P}_{\mathrm{c}1})$ up to order $y_{\mathrm{e}}^{-3}$ to show that the dependence on the details of the smoother transition only appear (through the parameter $X$) at this order.
Of course, the same steps and considerations can be followed to derive the expression of the coefficient $B_{2}$. We find that $B_{2}$ can be written as
$B_{2}$ $=A_{k}\sqrt{2}\,\mathrm{e}^{i\left[y_{\mathrm{e}}-\pi(1-q)/2\right]}y_{\mathrm{e}}^{1/2}X^{-1/6}\left[{\cal P}_{\mathrm{s}2}\sin\left(\zeta_{\mathrm{e}}+\frac{\pi}{4}\right)+{\cal P}_{\mathrm{c}2}\cos\left(\zeta_{\mathrm{e}}+\frac{\pi}{4}\right)\right],$ (6.19)
where $({\cal P}_{\mathrm{s}2},{\cal P}_{\mathrm{c}2})$ are given by
${\cal P}_{\mathrm{s}2}$ $=-\varepsilon\left[iT_{1/2-q}(y_{\mathrm{e}})+T^{\prime}_{1/2-q}(y_{\mathrm{e}})\right]\left[1-\frac{q(q-1)}{y_{\mathrm{e}}^{2}}\right]^{-1/4}A(\tau_{\mathrm{e}})$
$\quad+T_{1/2-q}(y_{\mathrm{e}})\left[1-\frac{q(q-1)}{y_{\mathrm{e}}^{2}}\right]^{1/4}D(\tau_{\mathrm{e}}),$ (6.20a)
${\cal P}_{\mathrm{c}2}$ $=\varepsilon\left[iT_{1/2-q}(y_{\mathrm{e}})+T^{\prime}_{1/2-q}(y_{\mathrm{e}})\right]\left[1-\frac{q(q-1)}{y_{\mathrm{e}}^{2}}\right]^{-1/4}B(\tau_{\mathrm{e}})$
$\quad+T_{1/2-q}(y_{\mathrm{e}})\left[1-\frac{q(q-1)}{y_{\mathrm{e}}^{2}}\right]^{1/4}C(\tau_{\mathrm{e}}).$ (6.20b)
In the limit $y_{\mathrm{e}}\gg 1$, the quantities $({\cal P}_{\mathrm{s}2},{\cal P}_{\mathrm{c}2})$ can be expressed as
${\cal P}_{\mathrm{s}2}$ $=-i\varepsilon+\frac{\varepsilon q(q-1)}{2y_{\mathrm{e}}}+\frac{i\varepsilon q^{2}(q-1)^{2}}{8y_{\mathrm{e}}^{2}}$
$\quad-\frac{1}{48y_{\mathrm{e}}^{3}}\left\{\left[12q-14q^{2}+q^{3}(3+q-3q^{2}+q^{3})\right]\varepsilon+7X\right\}+{\cal O}\left(\frac{1}{y_{\mathrm{e}}^{4}}\right),$ (6.21a)
${\cal P}_{\mathrm{c}2}$ $=1+\frac{iq(q-1)}{2y_{\mathrm{e}}}-\frac{q^{2}(q-1)^{2}}{8y_{\mathrm{e}}^{2}}$
$\quad+\frac{i}{48y_{\mathrm{e}}^{3}}\left[12q-10q^{2}-q^{3}(3+q-3q^{2}+q^{3})+5\varepsilon X\right]+{\cal O}\left(\frac{1}{y_{\mathrm{e}}^{4}}\right).$ (6.21b)
Thus, we have completely determined the coefficients $(B_{1},B_{2})$ during the transition. The next step consists of matching the solution for the recsaled mode function $\mu_{k}(\eta)$ during the transition to the solution during the radiation-dominated epoch. Such a matching would lead to explicit expressions for the coefficients $(\alpha^{\mathrm{r}}_{k},\beta^{\mathrm{r}}_{k})$. We shall now turn to the calculation.
6.2 Evolution and PS after the smoother transition
During the radiation dominated era, the rescaled mode function $\mu_{k}(\eta)$ is given by Eq. (2.14) with the functions $m_{k}(\eta)$ and $n_{k}(\eta)$ being given by Eq. (2.16). The coefficients $(\alpha^{\mathrm{r}}_{k},\beta^{\mathrm{r}}_{k})$ are obtained by matching the mode function and its derivative during the smooth linear transition [cf. Eq. (6.3)] to the mode function and its time derivative during the radiation dominated era. The matching is to be carried out at the end of the period of the transition, viz. at $\eta_{\mathrm{tr}}=\eta_{\mathrm{e}}+\Delta\eta$. On carrying out the matching, we find that the coefficients $(\alpha^{\mathrm{r}}_{k},\beta^{\mathrm{r}}_{k})$ can be written as
$\begin{pmatrix}\alpha^{\mathrm{r}}_{k}\\ \beta^{\mathrm{r}}_{k}\end{pmatrix}=\frac{1}{W_{k}^{\mathrm{r}}}\begin{pmatrix}T_{11}&T_{12}\\ T_{21}&T_{22}\end{pmatrix}\begin{pmatrix}B_{1}\\ B_{2}\end{pmatrix}$, (6.22)
where the matrix elements $(T_{11},T_{12},T_{21},T_{22})$ are given by
$T_{11}$ $=n_{k}^{\prime}(\eta_{\mathrm{tr}})i_{k}(\eta_{\mathrm{tr}})-n_{k}(\eta_{\mathrm{tr}})\frac{{\mathrm{d}}\tau}{{\mathrm{d}}\eta}\frac{{\mathrm{d}}i_{k}}{{\mathrm{d}}\tau}\biggl|_{\eta_{\mathrm{tr}}},$ (6.23a)
$T_{12}$ $=n_{k}^{\prime}(\eta_{\mathrm{tr}})j_{k}(\eta_{\mathrm{tr}})-n_{k}(\eta_{\mathrm{tr}})\frac{{\mathrm{d}}\tau}{{\mathrm{d}}\eta}\frac{{\mathrm{d}}j_{k}}{{\mathrm{d}}\tau}\biggl|_{\eta_{\mathrm{tr}}},$ (6.23b)
$T_{21}$ $=-m_{k}^{\prime}(\eta_{\mathrm{tr}})i_{k}(\eta_{\mathrm{tr}})+m_{k}(\eta_{\mathrm{tr}})\frac{{\mathrm{d}}\tau}{{\mathrm{d}}\eta}\frac{{\mathrm{d}}i_{k}}{{\mathrm{d}}\tau}\biggl|_{\eta_{\mathrm{tr}}},$ (6.23c)
$T_{22}$ $=-m_{k}^{\prime}(\eta_{\mathrm{tr}})j_{k}(\eta_{\mathrm{tr}})+m_{k}(\eta_{\mathrm{tr}})\frac{{\mathrm{d}}\tau}{{\mathrm{d}}\eta}\frac{{\mathrm{d}}j_{k}}{{\mathrm{d}}\tau}\biggl|_{\eta_{\mathrm{tr}}},$ (6.23d)
and $W_{k}^{\mathrm{r}}=4ik/\pi$ is the Wronskian of the functions $m_{k}(\eta)$ and $n_{k}(\eta)$ [see Eq. (3.3) and the comment following it]. To proceed further, we make use of the results $\mathrm{d}n_{k}/\mathrm{d}\eta=ikn_{k}$ and $\mathrm{d}i_{k}/\mathrm{d}\eta=(\mathrm{d}\tau/\mathrm{d}\eta)\mathrm{d}i_{k}/\mathrm{d}\tau=(\mathrm{d}\rho/\mathrm{d}\eta)\mathrm{d}i_{k}/\mathrm{d}\tau$. Together with Eq. (5.15), these lead to the following expression for $T_{11}$:
$T_{11}=ik\,n_{k}(\eta_{\mathrm{tr}})\left[i_{k}(\eta_{\mathrm{tr}})-\frac{iX^{1/3}}{y_{\mathrm{e}}}\frac{\mathrm{d}i_{k}}{\mathrm{d}\tau}\biggl|_{\eta_{\mathrm{tr}}}\right].$ (6.24)
Note that the function $n_{k}(\eta)$ has to be evaluated at
$k(\eta_{\mathrm{tr}}-\eta_{\mathrm{r}})=k(\eta_{\mathrm{e}}+\Delta\eta-\eta_{\mathrm{r}})=k\eta_{\mathrm{e}}\left(1-\frac{\Delta\eta}{|\eta_{\mathrm{e}}|}+\frac{\eta_{\mathrm{r}}}{|\eta_{\mathrm{e}}|}\right)=-y_{\mathrm{e}}Y,$ (6.25)
where $Y$ is given by [cf. Eq. (5.22a)]
$Y=-\left[q(q-1)\right]^{-1/3}\left(\frac{\Delta\eta}{|\eta_{\mathrm{e}}|}\right)^{1/3}\frac{\Gamma(1/3)}{3^{1/3}\Gamma(2/3)}\left(\frac{A_{1}+\sqrt{3}A_{2}}{A_{1}-\sqrt{3}A_{2}}\right).$ (6.26)
On the other hand, the Airy function appearing in $i_{k}(\eta)$ has to be calculated at the following value of the argument:
$\tau(\eta_{\mathrm{tr}})$ $\equiv\tau_{\mathrm{tr}}=-\alpha\left[k^{2}-U(\eta_{\mathrm{e}}+\Delta\eta)\right]=-\alpha k^{2}=-y_{\mathrm{e}}^{2}X^{-2/3}\,$ (6.27)
where we have made use of the expression (5.10) for $\alpha$. We find that the above expression for $T_{11}$ and, similar expressions for $T_{12}$, $T_{21}$ and $T_{22}$, can be written as
$T_{11}$ $=\sqrt{\frac{2}{\pi}}k\,\mathrm{e}^{-iy_{\mathrm{e}}Y}\left[\mathrm{Ai}(\tau_{\mathrm{tr}})-\frac{iX^{1/3}}{y_{\mathrm{e}}}\mathrm{Ai}^{\prime}(\tau_{\mathrm{tr}})\right],$ (6.28a)
$T_{12}$ $=\sqrt{\frac{2}{\pi}}k\,\mathrm{e}^{-iy_{\mathrm{e}}Y}\left[\mathrm{Bi}(\tau_{\mathrm{tr}})-\frac{iX^{1/3}}{y_{\mathrm{e}}}\mathrm{Bi}^{\prime}(\tau_{\mathrm{tr}})\right],$ (6.28b)
$T_{21}$ $=-\sqrt{\frac{2}{\pi}}k\,\mathrm{e}^{iy_{\mathrm{e}}Y}\left[\mathrm{Ai}(\tau_{\mathrm{tr}})+\frac{iX^{1/3}}{y_{\mathrm{e}}}\mathrm{Ai}^{\prime}(\tau_{\mathrm{tr}})\right],$ (6.28c)
$T_{22}$ $=-\sqrt{\frac{2}{\pi}}k\,\mathrm{e}^{iy_{\mathrm{e}}Y}\left[\mathrm{Bi}(\tau_{\mathrm{tr}})+\frac{iX^{1/3}}{y_{\mathrm{e}}}\mathrm{Bi}^{\prime}(\tau_{\mathrm{tr}})\right],$ (6.28d)
Finally, using the expressions (6.10) for the Airy function and its derivatives, we can write the matrix elements $(T_{11},T_{12},T_{21},T_{22})$ as follows:
$T_{11}$ $=k\frac{\sqrt{2}}{\pi}\mathrm{e}^{-iy_{\mathrm{e}}Y}y_{\mathrm{e}}^{-1/2}X^{1/6}\left[{\cal P}_{\mathrm{s11}}\sin\left(\zeta_{\mathrm{tr}}+\frac{\pi}{4}\right)+{\cal P}_{\mathrm{c11}}\cos\left(\zeta_{\mathrm{tr}}+\frac{\pi}{4}\right)\right],$ (6.29a)
$T_{12}$ $=k\frac{\sqrt{2}}{\pi}\mathrm{e}^{-iy_{\mathrm{e}}Y}y_{\mathrm{e}}^{-1/2}X^{1/6}\left[{\cal P}_{\mathrm{s12}}\sin\left(\zeta_{\mathrm{tr}}+\frac{\pi}{4}\right)+{\cal P}_{\mathrm{c12}}\cos\left(\zeta_{\mathrm{tr}}+\frac{\pi}{4}\right)\right],$ (6.29b)
$T_{21}$ $=-k\frac{\sqrt{2}}{\pi}\mathrm{e}^{iy\mathrm{e}Y}y_{\mathrm{e}}^{-1/2}X^{1/6}\left[{\cal P}_{\mathrm{s21}}\sin\left(\zeta_{\mathrm{tr}}+\frac{\pi}{4}\right)+{\cal P}_{\mathrm{c21}}\cos\left(\zeta_{\mathrm{tr}}+\frac{\pi}{4}\right)\right],$ (6.29c)
$T_{22}$ $=-k\frac{\sqrt{2}}{\pi}\mathrm{e}^{iy_{\mathrm{e}}Y}y_{\mathrm{e}}^{-1/2}X^{1/6}\left[{\cal P}_{\mathrm{s22}}\sin\left(\zeta_{\mathrm{tr}}+\frac{\pi}{4}\right)+{\cal P}_{\mathrm{c22}}\cos\left(\zeta_{\mathrm{tr}}+\frac{\pi}{4}\right)\right],$ (6.29d)
where the quantities $({\cal P}_{\mathrm{s11}},{\cal P}_{\mathrm{c11}},{\cal P}_{\mathrm{s12}},{\cal P}_{\mathrm{c12}},{\cal P}_{\mathrm{s21}},{\cal P}_{\mathrm{c21}},{\cal P}_{\mathrm{s22}},{\cal P}_{\mathrm{c22}})$ are given by
${\cal P}_{\mathrm{s11}}$ $\equiv A(\tau_{\mathrm{tr}})+iD(\tau_{\mathrm{tr}}),$ (6.30a)
${\cal P}_{\mathrm{c11}}$ $\equiv-B(\tau_{\mathrm{tr}})+iC(\tau_{\mathrm{tr}})$ (6.30b)
${\cal P}_{\mathrm{s12}}$ $\equiv B(\tau_{\mathrm{tr}})-iC(\tau_{\mathrm{tr}}),$ (6.30c)
${\cal P}_{\mathrm{c12}}$ $\equiv A(\tau_{\mathrm{tr}})+iD(\tau_{\mathrm{tr}}),$ (6.30d)
${\cal P}_{\mathrm{s21}}$ $\equiv A(\tau_{\mathrm{tr}})-iD(\tau_{\mathrm{tr}}),$ (6.30e)
${\cal P}_{\mathrm{c21}}$ $\equiv-B(\tau_{\mathrm{tr}})-iC(\tau_{\mathrm{tr}}),$ (6.30f)
${\cal P}_{\mathrm{s22}}$ $\equiv B(\tau_{\mathrm{tr}})+iC(\tau_{\mathrm{tr}}),$ (6.30g)
${\cal P}_{\mathrm{c22}}$ $\equiv A(\tau_{\mathrm{tr}})-iD(\tau_{\mathrm{tr}}).$ (6.30h)
The final step consists of making use of the above expressions for the matrix elements $(T_{11},T_{12},T_{21},T_{22})$ and the exact expressions for $(B_{1},B_{2})$ [cf. Eqs. (6.15) and (6.19)] in the matrix relation (6.22) to arrive at Bogoliubov coefficients during the epoch of radiation domination, viz. $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$. We find that $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ can be written as
$\alpha^{\mathrm{r}}_{k}$ $=-\frac{iA_{k}}{2}\mathrm{e}^{i[y_{\mathrm{e}}+(q-1)\pi/2]-iy_{\mathrm{e}}Y}\Biggl[\left({\cal P}_{\mathrm{s}11}{\cal P}_{\mathrm{c}1}+{\cal P}_{\mathrm{s}12}{\cal P}_{\mathrm{c}2}\right)\sin\left(\zeta_{\mathrm{tr}}+\frac{\pi}{4}\right)\cos\left(\zeta_{\mathrm{e}}+\frac{\pi}{4}\right)$
$\quad+\left({\cal P}_{\mathrm{s}11}{\cal P}_{\mathrm{s}1}+{\cal P}_{\mathrm{s}12}{\cal P}_{\mathrm{s}2}\right)\sin\left(\zeta_{\mathrm{tr}}+\frac{\pi}{4}\right)\sin\left(\zeta_{\mathrm{e}}+\frac{\pi}{4}\right)+\left({\cal P}_{\mathrm{c}11}{\cal P}_{\mathrm{c}1}+{\cal P}_{\mathrm{c}12}{\cal P}_{\mathrm{c}2}\right)\cos\left(\zeta_{\mathrm{tr}}+\frac{\pi}{4}\right)$
$\quad\times\cos\left(\zeta_{\mathrm{e}}+\frac{\pi}{4}\right)+\left({\cal P}_{\mathrm{c}11}{\cal P}_{\mathrm{s}1}+{\cal P}_{\mathrm{c}12}{\cal P}_{\mathrm{s}2}\right)\cos\left(\zeta_{\mathrm{tr}}+\frac{\pi}{4}\right)\sin\left(\zeta_{\mathrm{e}}+\frac{\pi}{4}\right)\Biggr],$ (6.31a)
$\beta^{\mathrm{r}}_{k}$ $=\frac{iA_{k}}{2}\mathrm{e}^{i[y_{\mathrm{e}}+(q-1)\pi/2]+iy_{\mathrm{e}}Y}\Biggl[\left({\cal P}_{\mathrm{s}21}{\cal P}_{\mathrm{c}1}+{\cal P}_{\mathrm{s}22}{\cal P}_{\mathrm{c}2}\right)\sin\left(\zeta_{\mathrm{tr}}+\frac{\pi}{4}\right)\cos\left(\zeta_{\mathrm{e}}+\frac{\pi}{4}\right)$
$\quad+\left({\cal P}_{\mathrm{s}21}{\cal P}_{\mathrm{s}1}+{\cal P}_{\mathrm{s}22}{\cal P}_{\mathrm{s}2}\right)\sin\left(\zeta_{\mathrm{tr}}+\frac{\pi}{4}\right)\sin\left(\zeta_{\mathrm{e}}+\frac{\pi}{4}\right)+\left({\cal P}_{\mathrm{c}21}{\cal P}_{\mathrm{c}1}+{\cal P}_{\mathrm{c}22}{\cal P}_{\mathrm{c}2}\right)\cos\left(\zeta_{\mathrm{tr}}+\frac{\pi}{4}\right)$
$\quad\times\cos\left(\zeta_{\mathrm{e}}+\frac{\pi}{4}\right)+\left({\cal P}_{\mathrm{c}21}{\cal P}_{\mathrm{s}1}+{\cal P}_{\mathrm{c}22}{\cal P}_{\mathrm{s}2}\right)\cos\left(\zeta_{\mathrm{tr}}+\frac{\pi}{4}\right)\sin\left(\zeta_{\mathrm{e}}+\frac{\pi}{4}\right)\Biggr].$ (6.31b)
Note that $\zeta_{\mathrm{tr}}$ and $\zeta_{\mathrm{e}}$ can be expressed in terms of $y_{\mathrm{e}}$ and $X$ as follows:
$\zeta_{\mathrm{tr}}$ $\equiv\frac{2}{3}(-\tau_{\mathrm{tr}})^{3/2}=\frac{2}{3}\frac{y_{\mathrm{e}}^{3}}{X},$ (6.32a)
$\zeta_{\mathrm{e}}$ $\equiv\frac{2}{3}(-\tau_{\mathrm{e}})^{3/2}=\frac{2}{3}\frac{y_{\mathrm{e}}^{3}}{X}\left[1-\frac{q(q-1)}{y_{\mathrm{e}}^{2}}\right]^{3/2}.$ (6.32b)
We should emphasize that the above expressions for the coefficients $(\alpha^{\mathrm{r}}_{k},\beta^{\mathrm{r}}_{k})$ are exact. In the limit $y_{\mathrm{e}}\gg 1$, it is easy to expand these expressions in inverse powers of $y_{\mathrm{e}}$. We find that, at the leading order, the coefficients $(\alpha^{\mathrm{r}}_{k},\beta^{\mathrm{r}}_{k})$ are given by
$\alpha_{k}^{\mathrm{r}}$ $\simeq-iA_{k}\mathrm{e}^{i[y_{\mathrm{e}}+(q-1)\pi/2]-iy_{\mathrm{e}}Y}\mathrm{e}^{i(\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}})},$ (6.33a)
$\beta_{k}^{\mathrm{r}}$ $\simeq\frac{q(q-1)A_{k}}{4y_{\mathrm{e}}^{3}}\mathrm{e}^{i[y_{\mathrm{e}}+\pi(q-1)\pi/2]+iy_{\mathrm{e}}Y}\mathrm{e}^{-i(\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}})}\left[1-\mathrm{e}^{i(\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}})}\frac{iX}{q(q-1)}\sin\left(\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}}\right)\right].$ (6.33b)
We should highlight the fact that the coefficient $\beta^{\mathrm{r}}_{k}$ behaves as $y_{\mathrm{e}}^{-3}$ in the limit $y_{\mathrm{e}}\gg 1$. In other words, $\beta_{k}^{\mathrm{r}}\propto k^{-3}$ in the limit $k\gg k_{\mathrm{e}}$. This should be contrasted with the result we had arrived at earlier in the case of instantaneous transition from inflation to the epoch of radiation domination. Recall that, in the case of instantaneous transition we had obtained that $\beta^{\mathrm{r}}_{k}\propto k^{-2}$ in the limit $k\gg k_{\mathrm{e}}$ [cf. Eq. (3.4b)]. This implies that the smoother transition has modified the scale dependence of the coefficient $\beta_{k}^{\mathrm{r}}$. Such a modification will, in turn change the behavior of the PS at large $k$. It should be clear that the behavior $\beta_{k}^{\mathrm{r}}\propto k^{-2}$ at large $k$ is an artifact of the instantaneous transition from inflation to radiation domination.
Let us now consider the limit of instantaneous transition. Evidently, this corresponds to the limit $\Delta\eta\to 0$ or, equivalently, $X\to\infty$ [cf. Eq. (5.9)]. Since we are also focusing on the limit $y_{\mathrm{e}}\gg 1$, we have [cf. Eq. (6.32)]
$\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}}=\frac{2}{3}\frac{y_{\mathrm{e}}^{3}}{X}\left\{\left[1-\frac{q(q-1)}{y_{\mathrm{e}}^{2}}\right]^{3/2}-1\right\}\simeq-\frac{q(q-1)}{X}y_{\mathrm{e}}\,$, (6.34)
and, hence, it follows that
$\frac{X}{q(q-1)}\sin(\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}})\simeq-y_{\mathrm{e}}.$ (6.35)
Moreover, in the limit $\Delta\eta\to 0$, on using the expressions (5.24), we obtain that
$\frac{A_{1}+\sqrt{3}A_{2}}{A_{1}-\sqrt{3}A_{2}}\simeq\frac{3^{1/3}\Gamma(2/3)}{q\Gamma(1/3)}X^{1/3},$ (6.36)
from which we can also deduce that, in this limit [cf. Eq. (6.26)]
$Y\simeq-q.$ (6.37)
For simplicity, let us restrict ourselves to the case of de Sitter inflation wherein q=1q=-1. In such a case, in the limits, $y_{\mathrm{e}}\gg 1$ and $\Delta\eta\to 0$, the expressions (6.33) for $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ reduce to
$\alpha^{\mathrm{r}}_{k}$ $\simeq iA_{k}\mathrm{e}^{2iy_{\mathrm{e}}},$ (6.38a)
$\beta^{\mathrm{r}}_{k}$ $\simeq\frac{-iA_{k}}{2y_{\mathrm{e}}^{2}}.$ (6.38b)
These are indeed the $y_{\mathrm{e}}\gg 1$ limits of the expressions (3.4) for $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ we obtained originally in the case of the instantaneous transition from de Sitter inflation to the epoch of radiation domination.
Figure 7: The actual and the regularized PS of PGWs, viz. $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ and $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$, evaluated during the early stages of the epoch of radiation epoch at the conformal time $\eta=2|\eta_{\mathrm{e}}|$ have been plotted (in red and blue) in the case of the smoother transition from inflation to radiation domination. As in Figs. 3, 4 and 5, we have set $g_{\ast,\mathrm{rh}}^{1/4}T_{\mathrm{rh}}=1.8\times 10^{16}\,\mathrm{GeV}$ in plotting this figure. Also, we have chosen the width of the smooth transition to be $\Delta\eta=2|\eta_{\mathrm{e}}|$. The effects of the smoother transition from inflation to radiation domination modify the PS over wave numbers such that $y_{\mathrm{e}}\gtrsim 1$. Hence, we have plotted the actual and regularized PS only over this domain. Clearly, the regularized PS oscillates about zero (over $y_{\mathrm{e}}\gg 1$), as in the case of the instantaneous transition, which we had illustrated in Fig. 4. However, note that, in the case of the smoother transition, the amplitude of the oscillations decreases with $k$ at large wave numbers.
Let us now discuss the effects of the smoother transition from inflation to radiation domination on the PS of PGWs. The actual and regularized PS of PGWs during the epoch of radiation domination, after the smoother transition from inflation, can be obtained by substituting the results (6.31) for $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ in Eqs. (2.3) and (4.14), respectively. In Fig. 7, we have plotted the actual and the regularized PS evaluated during the early stages of the epoch of radiation domination, after the smoother transition from inflation. In Fig. 8, we have compared the regularized PS in the cases of instantaneous and smoother transitions from inflation, when they are evaluated during the early stages of radiation domination. Two points should be clear from these figures. Firstly, in the case of the smoother transition, over wave numbers $k\gtrsim k_{\mathrm{e}}$, the regularized PS oscillates about zero as in the case of the instantaneous transition. Secondly, while the amplitude of the oscillations remain a constant in the case of the instantaneous transition, the amplitude decreases as $k^{-1}$ in the case of the smooth transition.
Figure 8: The regularized PS of PGWs $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$, evaluated during the early stages of the epoch of radiation domination at the conformal time $\eta=2|\eta_{\mathrm{e}}|$, have been plotted (in red and blue) in the cases of instantaneous and smoother transitions from inflation to radiation domination. The nature of the transition from inflation to radiation domination affects the PS only over wave numbers such that $y_{\mathrm{e}}\gtrsim 1$. Hence, as in the previous figure, we have illustrated the PS only over this domain. Note that, while the regularized PS has a constant amplitude over small scales (such that over $y_{\mathrm{e}}\gg 1$) in the case of the instantaneous transition, clearly, the amplitude of the oscillations decreases as $k^{-1}$ in the case of the smooth linear transition.
6.3 PS during matter domination, after the smoother transition
In this section, we shall discuss the behavior of PS of PGWs during the epoch of matter domination, after the smooth linear transition from inflation to the epoch of radiation domination. Earlier, we had expressed the actual and the regularized PS of PGWS during matter domination in terms of the Bogoliubov coefficients $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$ [see Eqs. (3.2) and (4.2)]. Recall that, these coefficients were, in turn, dependent on the Bogoliubov coefficients $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ during the epoch of radiation domination [see Eqs. (3.21)]. In the case of the smooth linear transition from inflation to radiation domination, we can substitute $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ from Eqs. (6.31) in Eqs. (3.21) to arrive at the coefficients $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$. The corresponding actual and regularized PS of PGWs during the epoch of matter domination can be arrived at by substituting these expressions for $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$ in Eqs. (3.2) and (4.2). In Fig. 9, we have plotted the actual and regularized PS of PGWs evaluated today. Evidently, the process regularization as well as the smoother transition from inflation to radiation domination affects the PS only over wave numbers such that $k\gtrsim k_{\mathrm{e}}$. Note that the process of regularization truncates the rise in the PS over $k\gtrsim k_{\mathrm{e}}$, as we have already discussed. The regularized PS oscillates about zero over this domain. The smooth linear transition leads to a suppression in the amplitude of the oscillations in the regularized PS (as $k^{-1}$) at such high wave numbers.
Figure 9: The actual and the regularized PS of PGWs, viz. $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ and $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$, evaluated today, i.e. during the late stages of matter domination at the conformal time $\eta_{0}$, have been plotted (in red and blue), assuming a smooth linear transition from de Sitter inflation to the epoch of radiation domination. In plotting the figure, we have assumed that the transition from radiation to matter domination is instantaneous. As in Figs. 3 and 5, the vertical lines indicate the frequencies associated with the wave numbers $(k_{0}/a_{0},k_{\mathrm{eq}}/a_{0},k_{\mathrm{e}}/a_{0})$. We have illustrated the PS only until the frequency of $f\simeq 10^{11}\,\mathrm{Hz}$, since, at higher frequencies, the numerical errors in plotting the regularized PS become significant. Note that the process of regularization does not alter the PS over wave numbers $k\lesssim k_{\mathrm{e}}$. However, as in Fig. 5, it is clear that regularization truncates the $k^{2}$ rise in the PS over $k\gtrsim k_{\mathrm{e}}$. In addition, the smooth linear transition from inflation to radiation domination suppresses the amplitude of the oscillations in the PS over $k\gtrsim k_{\mathrm{e}}$ as $k^{-1}$.
In the remainder of this section, we shall analytically arrive at the behavior of the PS during the epoch of matter domination on different scales. For simplicity, we shall assume de Sitter inflation corresponding to q=1q=-1. Let us first focus on wave numbers such that $y_{\mathrm{e}}\gg 1$. In such a limit, since $z\gg 1$ as well, the PS in Eq. (3.2) simplifies to be
[…]
Also, in the limit $x_{\mathrm{eq}}\gg 1$, we obtain that [cf. Eqs. (3.22)]
[…]
In arriving at these expressions, we have also made use of the fact that $1/x_{\mathrm{eq}}\gg 1/z_{0}$ and $1/y_{\mathrm{e}}^{3}\gg 1/(y_{\mathrm{e}}^{2}z_{0})$ since, $1/y_{\mathrm{e}}\gg 1/z_{0}$. Moreover, in such a limit, we have $\left|\alpha_{k}^{\mathrm{m}}/A_{k}\right|\simeq\left|\alpha_{k}^{\mathrm{r}}/A_{k}\right|\simeq 1$. Using the above results, we can write the PS during matter domination as
[…]
Further, from Eqs. (6.33), in the limit $y_{\mathrm{e}}\gg 1$, we have
[…]
$\quad+\frac{X}{2}\cos\left(\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}}-2Yy_{\mathrm{e}}\right)\sin\left(\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}}\right)\Biggr],$ (6.42a)
[…]
$\quad+\frac{X}{2}\sin\left(\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}}-2Yy_{\mathrm{e}}\right)\sin\left(\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}}\right)\Biggr].$ (6.42b)
As a result, the PS in the limit $y_{\mathrm{e}}\gg 1$ can be written as
$\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ $=\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}y_{\mathrm{e}}^{2}\Biggl\{1-\frac{1}{4x_{\mathrm{eq}}^{2}}\cos(4x_{\mathrm{eq}}-2z)$
$\quad-\frac{1}{y_{\mathrm{e}}^{3}}\sin\left[2\left(x_{\mathrm{eq}}-Yy_{\mathrm{e}}-z+\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}}\right)\right]$
$\quad-\frac{X}{2y_{\mathrm{e}}^{3}}\cos\left(2x_{\mathrm{eq}}-2Yy_{\mathrm{e}}-2z+\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}}\right)\sin\left(\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}}\right)\biggr\}.$ (6.43)
Note that, if we now consider the regularized PS, the first term will not be present. On comparing the last two terms with the second term, we find that, around $k\gtrsim k_{\mathrm{e}}$, the last two terms will dominate. These terms oscillate about zero as in the case of the regularized PS over this domain in the case of the instantaneous transition from inflation to radiation domination [see Eq. (4.20)]. But, because of the smoother transition from inflation to the epoch of radiation domination, the amplitude of the oscillations is suppressed as $k^{-1}$. However, interestingly, we find that the second term eventually becomes dominant in the regularized PS over $k\gtrsim k_{\mathrm{e}}^{3}/k_{\mathrm{eq}}^{2}$. Therefore, at ultra-high wave numbers, we find that the regularized PS behaves as
$\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ $=-\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\frac{y_{\mathrm{e}}^{2}}{4x_{\mathrm{eq}}^{2}}\cos(4x_{\mathrm{eq}}-2z).$ (6.44)
In other words, at ultra-high wave numbers, the amplitude of the oscillations in the regularized PS ceases to decrease as $k^{-1}$ and settles down to a constant value.
Let us now turn to the limit $y_{\mathrm{e}}\ll 1$. These correspond to wave numbers which are outside the Hubble radius at the end of inflation. In this regime, we find that the Bogoliubov coefficients $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ can be expanded as
$\alpha_{k}^{\mathrm{r}}$ $\simeq\mathcal{F}\frac{iA_{k}}{y_{\mathrm{e}}^{2}}+i\mathcal{G}\frac{iA_{k}}{y_{\mathrm{e}}}+\mathcal{O}(y_{\mathrm{e}}^{0}),$ (6.45a)
$\beta_{k}^{\mathrm{r}}$ $\simeq\mathcal{F}\frac{iA_{k}}{y_{\mathrm{e}}^{2}}-i\mathcal{G}\frac{iA_{k}}{y_{\mathrm{e}}}+\mathcal{O}(y_{\mathrm{e}}^{0}),$ (6.45b)
where $\mathcal{F}$ and $\mathcal{G}$ are dimensionless real functions which depend on the parameters of the smooth transition. They are given by
$\mathcal{F}$ $=-\frac{\pi}{6\Gamma(1/3)}\bigg[3^{7/6}\mathrm{Ai}(\tau_{\mathrm{e}})+3^{2/3}\mathrm{Bi}(\tau_{\mathrm{e}})+3^{7/6}X^{1/3}\mathrm{Ai}^{\prime}(\tau_{\mathrm{e}})+3^{2/3}X^{1/3}\mathrm{Bi}^{\prime}(\tau_{\mathrm{e}})\bigg],$ (6.46a)
$\mathcal{G}$ $=-\frac{1}{4\times 3^{1/3}}\bigg\{\left[-3^{2/3}\mathrm{Ai}^{\prime}(\tau_{\mathrm{e}})+3^{1/6}\mathrm{Bi}^{\prime}(\tau_{\mathrm{e}})-3^{2/3}X^{-1/3}\mathrm{Ai}(\tau_{\mathrm{e}})+3^{1/6}X^{-1/3}\mathrm{Bi}(\tau_{\mathrm{e}})\right]\Gamma\left(\frac{1}{3}\right)$
$\quad+\left[3\mathrm{Ai}(\tau_{\mathrm{e}})+3^{1/2}\mathrm{Bi}(\tau_{\mathrm{e}})+3X^{1/3}\mathrm{Ai}^{\prime}(\tau_{\mathrm{e}})+3^{1/2}X^{1/3}\mathrm{Bi}^{\prime}(\tau_{\mathrm{e}})\right]Y\,\Gamma\left(\frac{2}{3}\right)\bigg\},$ (6.46b)
where $\tau_{\mathrm{e}}=\tau(\eta_{\mathrm{e}})\simeq 2/X^{2/3}$ [cf. Eq. (6.9)].
We can now distinguish three different cases. The first corresponds to $y_{\mathrm{e}}\ll 1$, $x_{\mathrm{eq}}\gg 1$ and $z_{0}\gg 1$. These limits correspond to wave numbers $k_{\mathrm{eq}}\lesssim k\lesssim k_{\mathrm{e}}$, i.e. wave numbers which have re-entered the Hubble radius before the time of radiation-matter equality. In this case, the relations between the coefficients $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$ and $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ are given by Eqs. (3.23). On using the fact that $z\gg 1$, we find that the PS during matter domination, as given by Eq. (3.2), reduces to the following form:
[…]
The expressions (6.45) for $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ imply that, at leading order in the limit $y_{\mathrm{e}}\ll 1$,
$\left|\frac{\alpha_{k}^{\mathrm{r}}}{A_{k}}\right|^{2}=\left|\frac{\beta_{k}^{\mathrm{r}}}{A_{k}}\right|^{2}=\frac{\alpha_{k}^{\mathrm{r}}\beta_{k}^{\mathrm{r}*}}{|A_{k}|^{2}}\simeq\frac{{\cal F}^{2}}{y_{\mathrm{e}}^{4}}.$ (6.48)
It then follows that the above PS simplifies to
$\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)\simeq\frac{8\mathcal{F}^{2}}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\frac{1}{y_{\mathrm{e}}^{2}}\sin^{2}(x_{\rm eq}-z).$ (6.49)
The second case corresponds to $y_{\mathrm{e}}\ll 1$, $x_{\mathrm{eq}}\ll 1$ and $z_{0}\gg 1$. These limits correspond to the wave numbers $k_{0}\lesssim k\lesssim k_{\mathrm{eq}}$, which re-enter the Hubble radius during the epoch of matter domination. Recall that, in the domain $x_{\mathrm{eq}}\ll 1$, the relation between the coefficients $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$ and $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ are given by Eqs. (3.27). On making use of the expressions (6.45) for $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$, we obtain that
$\alpha_{k}^{\mathrm{m}}\simeq\beta_{k}^{\mathrm{m}}\simeq\frac{2iA_{k}}{8x_{\mathrm{eq}}^{2}y_{\mathrm{e}}^{2}}\left(-{\cal G}y_{\mathrm{e}}+3x_{\mathrm{eq}}{\cal F}\right),$ (6.50)
where we have retained the leading and next-to-leading orders in Eqs. (3.27). If we now make use of these results in Eq. (3.2) and consider the limit $z\gg 1$, we obtain the PS during the epoch of matter domination to be
[…]
$\simeq\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}y_{\mathrm{e}}^{2}2\,\left|\frac{\beta_{k}^{\mathrm{m}}}{A_{k}}\right|^{2}\left[1+\cos(2z)\right]$
$\simeq\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}y_{\mathrm{e}}^{2}\frac{1}{4x_{\mathrm{eq}}^{4}y_{\mathrm{e}}^{4}}\left(-{\cal G}y_{\mathrm{e}}+3x_{\mathrm{eq}}{\cal F}\right)^{2}\cos^{2}z$
$\simeq\frac{9{\cal F}^{2}}{2\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\frac{1}{x_{\mathrm{eq}}^{2}y_{\mathrm{e}}^{2}}\cos^{2}z,$ (6.51)
where in the last step we have used the fact that $x_{\mathrm{eq}}\gg y_{\mathrm{e}}$.
The third and last case corresponds to $y_{\mathrm{e}}\ll 1$, $x_{\mathrm{eq}}\ll 1$ and $z_{0}\ll 1$. These limit correspond to wave numbers $k\lesssim k_{0}$, which are yet to enter the Hubble radius. In this domain, the PS during the epoch of matter domination can be written as
$\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ $\simeq\frac{2}{\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}y_{\mathrm{e}}^{2}\left|\frac{\alpha_{k}^{\mathrm{m}}}{A_{k}}\right|^{2}\biggl[2\left(1+\frac{1}{z^{2}}\right)+2\left(1-\frac{1}{z^{2}}\right)\cos(2z)-\frac{4}{z}\sin(2z)\biggr]$
$\simeq\frac{1}{18\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\frac{z^{4}}{x_{\mathrm{eq}}^{4}y_{\mathrm{e}}^{2}}\left(-{\cal G}y_{\mathrm{e}}+3x_{\mathrm{eq}}{\cal F}\right)^{2},$ (6.52)
where, in the second expression, we have made use of the fact that $z\ll 1$. It follows that
$\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)\simeq\frac{{\cal F}^{2}}{2\pi^{2}}\left(\frac{H_{\mathrm{e}}}{M_{{}_{\mathrm{Pl}}}}\right)^{2}\left(\frac{a_{\mathrm{e}}}{a}\right)^{2}\frac{z^{4}}{x_{\mathrm{eq}}^{2}y_{\mathrm{e}}^{2}},$ (6.53)
which is scale-invariant.
Finally, let us point out that, in the limit of instantaneous transition from de Sitter inflation to the epoch of radiation domination, we have $X\rightarrow\infty,\penalty 10000\ Y\rightarrow-1,\penalty 10000\ (\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}})\rightarrow 2y_{\mathrm{e}}/X$. In such a limit, we find that $\mathcal{F}\simeq-1/2$ and $\mathcal{G}\simeq 2/(3X^{2})$ and the above expressions (6.49), (6.51), and (6.53) for the PS reduce to the expressions (3.26), (3.28) and (3.29) we had obtained earlier.
7 Summary and outlook
In this final section, after a quick summary of the results we have obtained, we shall provide a brief outlook.
7.1 Summary
The direct or indirect detection of PGWs remains the holy grail in cosmology. The detection of PGWs can help us determine the energy scale at which inflation occurred. As we discussed in the introduction, on the largest scales, the CMB carries imprints of PGWs. However, currently, there exists only an upper bound on the contributions of the primordial GWs to the anisotropies in the CMB [11, 12, 30]. The latest CMB observations constrain the tensor-to-scalar ratio to be $r<0.036$ [31], which corresponds to the upper bound on the Hubble scale during inflation to be $H_{{}_{\mathrm{I}}}/M_{{}_{\mathrm{Pl}}}\simeq 10^{-5}$. The recent detection of a SGWB by the PTAs has provided hope that direct observations of PGWs at small scales or high frequencies may be possible in the near future. In fact, as we outlined in the introduction, a variety of GW observatories have been proposed, which are expected to operate at frequencies even as high as a few hundred GHz. In such a situation, it becomes imperative to understand the behavior of the PS of GWs at very high frequencies.
In this work, we have examined the effects of regularization and smoothing of the transition from inflation to the epoch of radiation domination on the PS of GWs. We have argued that regularization is essential if we are to avoid the $k^{2}$ rise in the PS on small scales. We have adopted the method of adiabatic subtraction to regularize the PS. We have evolved the PGWs until the time of radiation-matter equality and today, and regularized the PS at these instances. Importantly, we find that regularization does not affect the standard predictions over the wave numbers $k\lesssim k_{\mathrm{e}}$. Moreover, we find that, in the case of instantaneous transition from de Sitter inflation to the epoch of radiation domination, the regularized PS of PGWs does not exhibit a rise over wave numbers such that $k\gtrsim k_{\mathrm{e}}$, i.e. over wave numbers that always remain inside the Hubble radius. Instead, the PS oscillates about zero with a constant amplitude in this domain. We also showed that the regularization of the PS is not sufficient. It is also necessary to smooth the transitions in the cosmological scenario. It is only the combination of regularization and smooth transitions that can lead to well-behaved correlation functions in real space. This is of course natural since realistic transitions are expected to be smooth and, in fact, they are expected to be infinitely continuous. We worked with a linear function to smooth the ‘effective potential’ $U(\eta)$ during the transition from inflation to the epoch of radiation domination. We obtained exact solutions to the mode functions describing the PGWs during the transition and evaluated the PS of GWs at the time of radiation-matter equality as well as today. We showed that, in contrast to the case of instantaneous transition, the smoother transition leads to a suppression in the regularized PS of PGWs over wave numbers $k\gtrsim k_{\mathrm{e}}$. Over such high wave numbers, while the regularized PS continues to oscillate about zero as in the case of the instantaneous transition, it does so with a decreasing amplitude. Further, we encountered an interesting behavior in the regularized PS evaluated today, after a smoother transition from inflation to radiation domination and an instantaneous transition from radiation to matter domination. In this scenario, we found that the suppression in the regularized PS ceases at ultra-high wave numbers such that $k\gg k_{\mathrm{e}}^{3}/k_{\mathrm{eq}}^{2}$ and the amplitude of the oscillations in the PS becomes constant thereafter. This behavior can be attributed to the assumption of an instantaneous transition from radiation to matter domination. The result clearly suggests that, unless all the intermediate transitions are smoothed, the PS of PGWs evaluated in a given epoch will carry imprints of previous, unrealistic, instantaneous transitions.
7.2 Outlook
There are a few directions in which the calculations we have carried out need to be extended. To begin with, we need to extend our analysis for transitions from inflation to radiation domination that are even more smooth. In fact, we have to work with $U(\eta)$ that are infinitely continuous. However, it is not easy to obtain solutions for the scalar factor $a(\eta)$ and the rescaled mode functions $\mu_{k}(\eta)$ in such situations. We may have to consider some approximations to determine the Bogoliubov coefficients $(\alpha_{k},\beta_{k})$ in these situations. Secondly, we need to introduce a phase of reheating between (inflation and the epoch of radiation domination) and evolve the PGWs across the phase of reheating as well as the epochs of radiation and matter domination, and evaluate the regularized PS of PGWs today. These are relatively straightforward exercises and can be implemented easily. Thirdly, we need to smooth the different transitions from the end of inflation until today and determine the corresponding imprints on the PS of PGWs. It seems clear that the suppression in the PS of PGWs over $k\gtrsim k_{\mathrm{e}}$ arises due to the smoothing of the transition from inflation to the epoch that succeeds it. The smoothing of the later transitions can be expected to lead to some features around wave numbers that re-enter the Hubble radius at the time of these transitions, say, $k_{\mathrm{re}}$, corresponding to the transition from the phase of reheating to the epoch of radiation domination, and $k_{\mathrm{eq}}$ corresponding to the time of transition from the epoch of radiation domination to that of matter domination. Fourthly, we have seen that the smoother transition from inflation to the epoch that follows will lead to a suppression in the regularized PS of PGWs over the wave numbers $k\gtrsim k_{\mathrm{e}}$. In the simplest of scenarios, a smooth transition from inflation to reheating and eventually to the epoch of radiation domination is achieved by coupling the inflation to radiation with the help of a constant coupling parameter. The manner in which the regularized PS of PGWs will be suppressed at high wave numbers can be expected to depend on the details of the transition. As is well known, the nature of the epoch of reheating will depend on the behavior of the inflationary potential near the minimum. It seems worthwhile to systematically examine the relation between the form of the suppression in the regularized PS and the behavior of the inflationary potential near its minimum and the coupling of the inflation to radiation. We are currently working on these problems.
Acknowledgements
The authors wish to thank Debika Chowdhury, Fabio Finelli, John Giblin, Vincent Vennin and Masahide Yamaguchi for discussions. AH wishes to thank Indian Institute of Technology (IIT) Madras, Chennai, India, for supporting a visit to Institut d’Astrophysique de Paris (IAP), France, through the International Immersion Experience Program. AP, JM and LS would like to thank the Indo-French Centre for the Promotion of Advanced Research (IFCPAR/CEFIPRA), New Delhi, India, for support of the proposal 6704-4 titled ‘Testing flavors of the early universe beyond vanilla models with cosmological observations’ under the Collaborative Scientific Research Programme. LS would also like to thank IAP and Fédération de Recherche Interactions Fondamentales for supporting a visit to IAP, where part of this work was completed. The authors also thank IAP and IIT Madras for hospitality.
Appendix A Equation of motion of PGWs
In this appendix, we shall derive the equation of motion governing PGWs. Though this is a standard result, we shall present the derivation for the sake of completeness. Also, the derivation helps us to set the notation.
Let us expand the metric tensor as
$g_{\mu\nu}=\bar{g}_{\mu\nu}+\tilde{h}_{\mu\nu}+j_{\mu\nu},$ (A.1)
where $\bar{g}_{\mu\nu}$ is the background metric, $\tilde{h}_{\mu\nu}$ is the first order perturbation in the metric, and $j_{\mu\nu}$ refers to the second order perturbations in the metric, i.e. it is of $\mathcal{O}(\tilde{h}^{2})$. Using this perturbative expansion, we can express the Ricci tensor as follows:
$R_{\mu\nu}=\bar{R}_{\mu\nu}+R_{\mu\nu}^{(1)}(\bar{g},\tilde{h})+R_{\mu\nu}^{(2)}(\bar{g},\tilde{h}^{2})+R_{\mu\nu}^{(2)}(\bar{g},j),$ (A.2)
where the first term is the Ricci tensor associated with the background, the second and third terms are linear and quadratic in $\tilde{h}_{\mu\nu}$, respectively, and the last term, as we have indicated, is the contribution from $j_{\mu\nu}$.
To arrive at the equation of motion of GWs, we begin with the following expression for the Ricci tensor $R_{\mu\nu}^{(1)}$ at the first order in the perturbations (in this context, see Ref. [115], Sec. 35.13):
$R_{\mu\nu}^{(1)}=\frac{1}{2}\left(-\nabla_{\nu}\nabla_{\mu}\tilde{h}-\nabla^{\alpha}\nabla_{\alpha}\tilde{h}_{\mu\nu}+\nabla^{\alpha}\nabla_{\nu}\tilde{h}_{\alpha\mu}+\nabla^{\alpha}\nabla_{\mu}\tilde{h}_{\alpha\nu}\right),$ (A.3)
where the covariant derivatives are to be taken with respect to the background metric $\bar{g}_{\mu\nu}$. Let us now assume that $\bar{g}_{\mu\nu}$ denotes the line-element (1.2) expressed in terms of the cosmic time $t$. The time-time-component of $R_{\mu\nu}^{(1)}$ is given by
$R_{tt}^{(1)}=\frac{1}{2}\left(-\nabla_{t}\nabla_{t}\tilde{h}-\nabla^{\alpha}\nabla_{\alpha}\tilde{h}_{tt}+\nabla^{\alpha}\nabla_{t}\tilde{h}_{\alpha t}+\nabla^{\alpha}\nabla_{t}\tilde{h}_{\alpha t}\right).$ (A.4)
Note that, in the case of GWs, $\tilde{h}_{t\mu}=\tilde{h}^{t\mu}=0$ [cf. Eq. (2.1)]. Also, the transverse and traceless conditions (2.6) correspond to $\nabla_{\mu}\tilde{h}^{\mu\nu}=0$ and $\tilde{h}=\tilde{h}^{\mu}_{\mu}=0$. On using these conditions, the different terms of $R_{tt}^{(1)}$ can be calculated to be
$\nabla_{t}\nabla_{t}\tilde{h}$ $=0,$ (A.5a)
$\nabla^{\alpha}\nabla_{\alpha}\tilde{h}_{tt}$ $=0,$ (A.5b)
$\nabla^{\alpha}\nabla_{t}\tilde{h}_{\alpha t}$ $=0,$ (A.5c)
which lead to
$R_{tt}^{(1)}=0.$ (A.6)
The time-space components of $R_{\mu\nu}^{(1)}$ are given by
$R_{ti}^{(1)}=\frac{1}{2}\left(-\nabla_{i}\nabla_{t}\tilde{h}-\nabla^{\alpha}\nabla_{\alpha}\tilde{h}_{ti}+\nabla^{\alpha}\nabla_{i}\tilde{h}_{\alpha t}+\nabla^{\alpha}\nabla_{t}\tilde{h}_{\alpha i}\right).$ (A.7)
The different terms of $R_{ti}^{(1)}$ can be evaluated to be
$\nabla_{i}\nabla_{t}\tilde{h}$ $=0,$ (A.8a)
$\nabla^{\alpha}\nabla_{\alpha}\tilde{h}_{ti}$ $=0,$ (A.8b)
$\nabla^{\alpha}\nabla_{i}\tilde{h}_{\alpha t}$ $=0,$ (A.8c)
$\nabla^{\alpha}\nabla_{t}\tilde{h}_{\alpha i}$ $=0,$ (A.8d)
leading to
$R_{ti}^{(1)}=0.$ (A.9)
Lastly, the space-space components of $R_{\mu\nu}^{(1)}$ are given by
$R_{ij}^{(1)}=\frac{1}{2}\left(-\nabla_{j}\nabla_{i}\tilde{h}-\nabla^{\alpha}\nabla_{\alpha}\tilde{h}_{ij}+\nabla^{\alpha}\nabla_{j}\tilde{h}_{\alpha i}+\nabla^{\alpha}\nabla_{i}\tilde{h}_{\alpha j}\right).$ (A.10)
The different terms of $R_{ij}^{(1)}$ can be calculated to be
$\nabla_{j}\nabla_{i}\tilde{h}$ $=0,$ (A.11a)
$\nabla^{\alpha}\nabla_{\alpha}\tilde{h}_{ij}$ $=-\ddot{\tilde{h}}_{ij}+H\dot{\tilde{h}}_{ij}+2\dot{H}\tilde{h}_{ij}+4H^{2}\tilde{h}_{ij}+\frac{1}{a^{2}}\partial^{2}\tilde{h}_{ij},$ (A.11b)
$\nabla^{\alpha}\nabla_{j}\tilde{h}_{\alpha i}$ $=\dot{H}\tilde{h}_{ij}+4H^{2}\tilde{h}_{ij},$ (A.11c)
$\nabla^{\alpha}\nabla_{i}\tilde{h}_{\alpha j}$ $=\dot{H}\tilde{h}_{ij}+4H^{2}\tilde{h}_{ij},$ (A.11d)
where, recall that, $H=\dot{a}/a$ is the Hubble parameter and $\partial^{2}=\delta^{ij}\partial_{i}\partial_{j}$. These results lead to
$R_{ij}^{(1)}=\frac{1}{2}\left(\ddot{\tilde{h}}_{ij}-H\dot{\tilde{h}}_{ij}+4H^{2}\tilde{h}_{ij}-\frac{1}{a^{2}}\partial^{2}\tilde{h}_{ij}\right).$ (A.12)
and, since $\tilde{h}_{ij}=a^{2}h_{ij}$ [cf. Eqs. (A.1) and (2.1)], this expression translates to
$R_{ij}^{(1)}=\frac{a^{2}}{2}\,\left(\ddot{h}_{ij}+3H\dot{h}_{ij}+4H^{2}h_{ij}+2\frac{\ddot{a}}{a}h_{ij}-\frac{1}{a^{2}}\partial^{2}h_{ij}\right).$ (A.13)
At the first order in the perturbations, the Einstein tensor is given by
$G_{\mu\nu}^{(1)}=R_{\mu\nu}^{(1)}-\frac{1}{2}g_{\mu\nu}^{(1)}R^{(0)}=R_{\mu\nu}^{(1)}-3a^{2}\left(\frac{\ddot{a}}{a}+H^{2}\right)h_{\mu\nu}.$ (A.14)
Since $h_{t\mu}=0$ and $R_{t\mu}^{(1)}=0$, we obtain that $G_{t\mu}^{(1)}=0$. At the same order, the spatial components of the Einstein tensor can be evaluated to be
$G_{ij}^{(1)}=\frac{a^{2}}{2}\left(\ddot{h}_{ij}+3H\dot{h}_{ij}-\frac{1}{a^{2}}\partial^{2}h_{ij}\right)-a^{2}\left(2\frac{\ddot{a}}{a}+H^{2}\right)h_{ij}.$ (A.15)
Recall that the stress-energy tensor associated with a perfect fluid is given by
$T_{\mu\nu}=(\rho+p)u_{\mu}u_{\nu}+pg_{\mu\nu},$ (A.16)
where $\rho$ and $p$ denote the energy density and pressure of the fluid and $u^{\mu}$ represents its four velocity, which satisfies the condition $g_{\mu\nu}u^{\mu}u^{\nu}=-1$. We should mention that $(\rho,p,u^{\mu})$ include the perturbations as well. According to the decomposition theorem, which is valid at the first order in perturbation theory, the GWs will be sourced only by the tensor component of the above stress-energy tensor (for a discussion in this context, see the textbook [96]). At the first order in the perturbations, the tensor component of the stress-energy tensor in the FLRW universe is given by
$T_{t\mu}^{(1)}$ $=0,$ (A.17a)
$T_{ij}^{(1)}$ $=\bar{p}a^{2}h_{ij},$ (A.17b)
where $\bar{p}$ is the background pressure that is driving the expansion. Note that, according to the zeroth order Friedmann equation,
$2\frac{\ddot{a}}{a}+H^{2}=-8\pi G\bar{p}.$ (A.18)
On taking into account the above results, the first order Einstein equation $G_{ij}^{(1)}=8\pi GT_{ij}^{(1)}$ leads to the following equation of motion governing the PGWs:
$\ddot{h}_{ij}+3H\dot{h}_{ij}-\frac{1}{a^{2}}\partial^{2}h_{ij}=0.$ (A.19)
In terms of the conformal time $\eta$, this equation reduces to
$h_{ij}^{\prime\prime}+2{\cal H}h_{ij}^{\prime}-\partial^{2}h_{ij}=0,$ (A.20)
where ${\cal H}=a^{\prime}/a=aH$ is the conformal Hubble parameter.
Appendix B Characteristic strain, sensitivity curves and the PS today
The PGWs will appear as a stochastic and isotropic background of GWs today. The characteristic strain $h_{\mathrm{c}}(f)$, the spectral density $S_{h}(f)$ and the PS $\mathcal{P}_{{}_{\mathrm{T}}}(f)$ are equivalent theoretical measures to describe a stochastic and isotropic GW background [97, 116, 98]. In this appendix, we shall discuss the relation between these quantities.
Let us decompose the stochastic GWs at a given spatial point, say, ${\bm{x}}=0$, as follows (in this regard, see Refs. [97, 116, 98]):
$h_{ij}(t)=\sum_{\lambda={(+,\times)}}\int_{-\infty}^{\infty}{\mathrm{d}}f\int{\mathrm{d}}\hat{\bm{n}}\,\varepsilon^{\lambda}_{ij}(\hat{\bm{n}})\tilde{h}^{\lambda}_{f}(\hat{\bm{n}})\mathrm{e}^{-2\pi ift},$ (B.1)
[…]
$\left\langle\tilde{h}_{f}^{\lambda}(\hat{\bm{n}})\tilde{h}_{f^{\prime}}^{\lambda^{\prime}\ast}(\hat{\bm{n}}^{\prime})\right\rangle=\delta^{(1)}(f-f^{\prime})\frac{1}{4\pi}\delta^{(2)}(\hat{\bm{n}},\hat{\bm{n}}^{\prime})\delta_{\lambda\lambda^{\prime}}\frac{1}{2}S_{h}(f),$ (B.2)
where the angular brackets on the left-hand side denote averaging over the random variables, and the quantity $S_{h}(f)$ is referred to as the spectral density of GWs. We should point out that the delta function $\delta^{(2)}(\hat{\bm{n}},\hat{\bm{n}}^{\prime})$ in the above expression implies that the stochastic background is isotropic. Note that the delta function satisfies the condition
$\int{\mathrm{d}}^{2}\hat{\bm{n}}\,\delta^{(2)}(\hat{\bm{n}},\hat{\bm{n}}^{\prime})=\int_{0}^{2\pi}{\mathrm{d}}\phi\,\delta^{(1)}(\phi-\phi^{\prime})\int_{0}^{\pi}{\mathrm{d}}\theta\sin\theta\,\delta^{(1)}(\cos\theta-\cos\theta^{\prime})=1.$ (B.3)
If we now substitute the decomposition (B.1) in Eq. (B.2), we obtain that
$\delta^{im}\delta^{jn}\langle h_{mn}(t)h_{ij}(t)\rangle=2\int_{-\infty}^{\infty}{\mathrm{d}}fS_{h}(f)=4\int_{0}^{\infty}{\mathrm{d}}(\ln f)fS_{h}(f),$ (B.4)
where we have made use of the fact that $S_{h}(f)=S_{h}(-f)$. It should be clear from the above relation that $S_{h}(f)$ has the same dimensions as $f^{-1}$. The so-called dimensionless characteristic strain $h_{\mathrm{c}}(f)$, which is a measure of the amplitude of GWs as a function of frequency, is defined through the relation [97, 98]
$\delta^{im}\delta^{jn}\langle h_{mn}(t)h_{ij}(t)\rangle=2\int_{0}^{\infty}{\mathrm{d}}(\ln f)h_{\mathrm{c}}^{2}(f)$ (B.5)
and, from the above two equations, we obtain that
$h_{\mathrm{c}}^{2}(f)=2fS_{h}(f).$ (B.6)
On comparing Eqs. (2.10) and (B.4), we arrive at the following relation between the PS of PGWs today, viz. $\mathcal{P}_{{}_{\mathrm{T}}}(f,\eta_{0})$, the spectral density $S_{h}(f)$ and the characteristic strain $h_{\mathrm{c}}(f)$:
$\mathcal{P}_{{}_{\mathrm{T}}}(f,\eta_{0})=4fS_{h}(f)=2h_{\mathrm{c}}^{2}(f).$ (B.7)
Often, the sensitivity curves of the GW observatories are indicated in terms of the characteristic strain $h_{\mathrm{c}}(f)$ (for related discussions, see Ref. [102] and the associated web-page; in this regard, also see Refs. [116, 103]). In Fig. 10, we have illustrated these sensitivity curves for the different GW observatories.
Figure 10: The sensitivity curves for the dimensionless characteristic strain $h_{\mathrm{c}}(f)$ have been illustrated for the different GW observatories [102, 103].
We have made use of the sensitivity curves for $h_{\mathrm{c}}(f)$ and the relation (B.7) above to arrive at the corresponding sensitivity curves for the PS of GWs. It is these sensitivity curves that we have included in Figs. 3, 5 and 9.